### v1

# GeneMamba: Efficient and Effective Foundation Model on Single Cell Data

Abstract
Transformer-based architectures have shown promise across various domains but struggle with computational inefficiency and scalability. To address these challenges, we introduce GeneMamba, a novel model designed specifically for single-cell data analysis. GeneMamba incorporates the BiMamba module to efficiently capture gene context information and employs biologically meaningful loss functions during training. The model enables scalable processing of over 50 million cells while significantly reducing computational costs. It delivers strong performance in multi-batch integration, cell type annotation, and gene pair correlation analysis. Furthermore, reconstruction experiments highlight GeneMamba’s explainability, establishing it as a robust foundation for advancing single-cell transcriptomics in biological and biomedical research.
1 Introduction
Single-cell RNA sequencing (scRNA-seq) has fundamentally transformed our ability to study cellular heterogeneity and dynamics by enabling high-resolution profiling of gene expression at the individual cell level (Korsunsky et al., 2019; Theodoris et al., 2023; Smith & Doe, 2024). This technology allows researchers to dissect complex cellular compositions within tissues, trace differentiation trajectories, and identify rare cell populations that remain undetectable with bulk RNA sequencing (Shen et al., 2023; Wen et al., 2023b; Nguyen & Tran, 2024). The rich transcriptomic insights provided by scRNA-seq have catalyzed significant advancements in developmental biology, disease modeling, and drug discovery (Gu & Dao, 2023; Yang et al., 2024; Li & Wang, 2024).
Despite these advantages, the computational analysis of scRNA-seq data presents formidable challenges due to its high dimensionality, inherent sparsity, and technical noise (Du et al., 2019; Zhao et al., 2023; Brown & Davis, 2024). Addressing these challenges requires robust computational models capable of capturing biologically meaningful patterns while efficiently handling the scale and variability of single-cell data. Transformer-based architectures, such as scBERT (Yang et al., 2022) and scGPT (Cui et al., 2024a), have emerged as powerful tools in this domain, demonstrating strong performance in tasks such as cell type classification, gene expression imputation, and differential expression analysis (Bian et al., 2024; Wen et al., 2023a; Garcia & Thompson, 2024).
However, transformers exhibit critical limitations when applied to scRNA-seq data. The quadratic complexity of their self-attention mechanism constrains their scalability for long sequences, which are characteristic of single-cell transcriptomes (Vaswani, 2017; Choromanski et al., 2020; Smith & Doe, 2024). Moreover, transformers often struggle to effectively capture long-range dependencies, which are essential for modeling gene regulatory interactions and cell state transitions (Dao et al., 2022; Chen et al., 2023). These limitations have driven the exploration of alternative architectures, among which state space models (SSMs) have emerged as promising solutions, offering improved efficiency and scalability for processing long sequences (Gu & Dao, 2023; Dao & Gu, 2024; Martinez & Hernandez, 2024).
In this study, we introduce GeneMamba, a novel state space model designed to efficiently train large-scale cell models on scRNA-seq data. SSM-based models have demonstrated competitive or superior performance compared to transformers while significantly reducing computational overhead (Shen et al., 2022; Gong et al., 2024a; Lee & Kim, 2024). Building upon this foundation, GeneMamba incorporates bidirectional computation, enhancing its ability to capture both upstream and downstream contextual dependencies, thereby improving its performance in tasks requiring global contextual awareness (Liu et al., 2024; Liang et al., 2024). We validate GeneMamba across a diverse set of applications, including multi-batch integration, cell type annotation, and gene-gene correlation analysis. Our experimental results demonstrate substantial improvements in computational efficiency and predictive performance compared to existing models (Hao et al., 2024; Sun et al., 2024; Johnson & Lee, 2024). Furthermore, to assess its scalability, we conduct a comprehensive performance analysis on various variations of GeneMamba (Appendix D), highlighting its advantages in real-world biological applications.
Our contributions are threefold:
1. We present GeneMamba, a scalable and efficient model designed for single-cell RNA sequencing data, which harnesses the strengths of the SSM architecture. GeneMamba’s architecture is tailored to tackle the challenges of high-dimensional and sparse scRNA-seq data, offering a robust and flexible framework for single-cell analysis.
2. We validate the effectiveness of GeneMamba in multiple downstream tasks, showcasing its potential to advance single-cell transcriptomics research. Extensive experiments highlight GeneMamba’s versatility and robustness, making it a valuable tool for the single-cell research community.
3. We demonstrate GeneMamba’s superior reconstruction ability compared to transformer-based models, providing insights into the interpretability and effectiveness of state space models.
2 Related Work
2.1 Discretization of Gene Expression
Discretizing gene expression levels into tokens is a crucial step in single-cell transcriptomics modeling. Existing methods employ various tokenization strategies, each with distinct advantages and limitations.
Bin-based discretization, as used by scBERT (Yang et al., 2022), scGPT (Cui et al., 2024a), and scMulan (Bian et al., 2024), groups expression values into predefined bins. This approach preserves absolute value distributions and simplifies sequence modeling, but may introduce information loss, particularly for genes with subtle but biologically significant expression differences. Additionally, binning can be sensitive to parameter selection, affecting downstream results.
Value projection(Szałata et al., 2024), adopted by scFoundation (Hao et al., 2024) and its backbone model xTrimoGene (Gong et al., 2024a), projects gene expression values into continuous embeddings rather than discrete categories. This method maintains full data resolution by applying a linear transformation to the gene expression vector, which is then combined with gene-specific embeddings. However, the use of continuous embeddings diverges from traditional tokenization strategies in NLP-based transformers, and its impact on model performance remains an open question.
Rank-based discretization, utilized by Geneformer (Theodoris et al., 2023), GeneCompass (Yang et al., 2024), tGPT(Shen et al., 2023) and LangCell (Zhao et al., 2024b), transforms gene expression values into ordinal rankings. This approach effectively captures relative expression levels and is more robust to batch effects and noise. Our method is based on rank-based discretization, as employed in Geneformer, which aligns naturally with biological processes such as regulatory interactions.
2.2 Model Architectures for Single-Cell Analysis
Transformer-based architectures, used by scBERT, scGPT, Geneformer, and scFoundation, have been successfully applied to single-cell data, leveraging the power of deep learning to model complex gene expression patterns. These models have shown promising results in various tasks, including cell-type annotation, batch integration, and multiomics analysis. However, transformer(Vaswani, 2017) has inherent limitations. The most significant one is their quadratic computational complexity with respect to sequence length, which makes them less feasible for long sequences typical of scRNA-seq data. This issue arises because the self-attention mechanism used in transformer requires computing attention scores for all pairs of tokens, leading to inefficiencies in handling long sequences. Additionally, transformer often struggles with capturing long-range dependencies in sequences, which is crucial for understanding gene regulatory networks and cell state transitions. Mamba(Gu & Dao, 2023) has been introduced to address the transformer efficiency problem. Bidirectional Mamba(Bi-Mamba) has recently been used to solve long-sequence problems in tasks such as feature extraction(Sun et al., 2024), sequential recommendation(Liu et al., 2024), and time series forecasting(Liang et al., 2024). Bi-Mamba is designed to efficiently process ultra-long sequences with linear computational complexity, offering a significant reduction in memory and computation requirements compared to transformer. By leveraging state-space dynamics, Bi-Mamba can capture long-range dependencies effectively, making it well-suited for modeling gene regulatory interactions and cell state transitions. Furthermore, Bi-Mamba’s bidirectional processing enables the simultaneous consideration of upstream and downstream contexts, enhancing its ability to model complex dependencies in single-cell data. This architecture represents a promising alternative to transformer-based methods, addressing their key limitations while maintaining high performance in various single-cell analysis tasks.
More related work is provied in Appendix C.
3 Methods
In this section, we introduce the framework of GeneMamba (Figure 1), the designed BiMamba module (Figure 2), and the pretraining objective (Section 3.3).
Figure 1: The GeneMamba architecture and its downstream task applications. The framework begins with the collection of training data (approximately 50M cells) from CELLXGENE, encompassing a diverse array of tissues and organs. After preprocessing, the data is prepared through a Gene Rank Module to transform single-cell data into input sequences. The GeneMamba module then captures the contextual information within each single cell. Once pretrained, the model and its embeddings are employed for various downstream tasks to evaluate the model’s performance.
Figure 2: The schematic overview of BiMamba Block. The BiMamba Block processes input sequences bidirectionally, capturing forward and reverse context through shared convolutional layers (Conv) and structured state machines (SSM). A gating mechanism integrates the outputs, followed by linear projection and nonlinearity layers, generating a context-aware representation for downstream tasks.
3.1 Data Processing
A significant challenge in modeling single-cell data is that the input data is not a plain token sequence; instead, it comprises both gene tokens and their corresponding expression values. To address this, we represent the input data as a gene expression matrix $M\in\mathbb{R}^{c\times g}$, where $c$ is the number of cells, and $g$ is the number of genes. In this matrix, rows correspond to cells, columns correspond to genes, and unexpressed genes have a zero expression value.
First, we preprocess the single-cell data using standard techniques (Appendix A). Next, we normalize the matrix to account for sequencing depth and gene-specific variation. Specifically, each element $M_{ij}$ (the expression value of gene $j$ in cell $i$) is first divided by the total expression of all genes in cell $i$. Then, we compute the median of the non-zero expression values for each gene $j$ across all cells using the t-digest algorithm for efficient computation. The final normalized expression value is given by:
$M_{ij}^{\text{norm}}=\frac{M_{ij}/\sum_{k=1}^{n}M_{ik}}{\text{t-digest}\{M_{kj}\mid M_{kj}>0\}}$
Finally, we rank the genes within each cell in descending order based on their normalized expression values. This ranking approach highlights genes that distinguish cell states while deprioritizing universally high-expression housekeeping genes, ensuring they are assigned lower ranks in downstream analyses.
$R_{i}=\operatorname{argsort}\left(-M_{ij}^{\text{norm}}\right),\quad\forall j$
3.2 Bi-Mamba Architecture
State Space Models (SSMs) provide a powerful framework for modeling sequences by utilizing a latent state that evolves over time. At each time step, the latent state $h_{t}$ is updated based on the previous state $h_{t-1}$, the current input $x_{t}$, and the system’s parameters. The dynamics of the system are described as:
$h_{t}=Ah_{t-1}+Bx_{t},\quad y_{t}=Ch_{t}$
where $A,B,$ and $C$ are matrices defining how the input and state interact. This formulation enables the model to represent long-range dependencies in sequences efficiently. The convolutional view reformulates the output $y$ as:
$y=x*K,\quad K=[CB,CAB,CA^{2}B,\dots]$
where $K$ is a sequence-length-dependent kernel capturing temporal relationships.
However, traditional SSMs are limited by their static nature; parameters like $A,B,C$ are constant, making them unable to adapt dynamically to the input content. To address this, the Mamba model introduces input-dependent dynamics.
In Mamba, the parameters $A,B,\Delta$, and $C$ are functions of the input, allowing the system to dynamically adjust to the sequence content. The updated equations are:
$h_{t}=f_{A}(x_{t})h_{t-1}+f_{B}(x_{t})x_{t},\quad y_{t}=f_{C}(x_{t})h_{t}$
where $f_{A},f_{B},$ and $f_{C}$ are learned transformations of the input $x_{t}$. By making these parameters input-dependent, Mamba enables selective propagation of relevant information while filtering out noise, significantly enhancing the model’s ability to capture complex sequence dynamics. Mamba is also computationally efficient, achieving linear scaling with sequence length.
While Mamba processes sequences in a unidirectional manner, many tasks require bidirectional context to fully capture dependencies. To address this limitation, we extend Mamba to a bidirectional version, called Bi-Mamba.
The Bi-Mamba model processes the input sequence in both forward and reverse directions to capture bidirectional contextual relationships. Given an input sequence $S=[s_{1},s_{2},\dots,s_{n}]$, the following steps outline the processing pipeline in Bi-Mamba:
1. Reversing the Input: A reversed version of the sequence, $S_{\text{rev}}=[s_{n},s_{n-1},\dots,s_{1}]$, is created. During the reversing and combining process, padding tokens are handled separately to prevent alignment artifacts. Only valid tokens (e.g., sequence content) are flipped, while padding remains static. This ensures that information flows in both directions during processing.
2. Parallel Processing: Both the original sequence $S$ and the reversed sequence $S_{\text{rev}}$ are processed independently using identical Mamba layers with shared weights. For each layer $l$, the outputs are:
$h_{t}^{(l)}=f_{A}^{(l)}(s_{t})h_{t-1}^{(l)}+f_{B}^{(l)}(s_{t})s_{t}$
for the original sequence, and:
$\tilde{h}_{t}^{(l)}=f_{A}^{(l)}(s_{t})\tilde{h}_{t-1}^{(l)}+f_{B}^{(l)}(s_{t})s_{t}$
for the reversed sequence $S_{\text{rev}}$, where $\tilde{h}_{t}^{(l)}$ represents the reversed latent state.
3. Combining the Outputs: The outputs from the forward and reversed passes are combined to form a unified representation. Instead of summing the outputs directly, Bi-Mamba introduces a learnable gating mechanism to balance the contributions of forward and backward information:
$z_{t}^{(l)}=\sigma(W^{(l)}[h_{t}^{(l)},\tilde{h}_{t}^{(l)}])$
$o_{t}^{(l)}=z_{t}^{(l)}\cdot h_{t}^{(l)}+(1-z_{t}^{(l)})\cdot\tilde{h}_{t}^{(l)}$
where $\sigma$ is a sigmoid function, $W^{(l)}$ is a learnable weight matrix, and $o_{t}^{(l)}$ is the combined output for token $t$ at layer $l$.
4. Stacking Layers: The Bi-Mamba process is repeated across multiple layers, with each layer refining the bidirectional representations:
$S^{(l+1)}=\text{Bi-Mamba}(S^{(l)}),\quad S^{(0)}=S$
The final sequence representation $S^{(L)}$ encodes rich bidirectional dependencies, making it suitable for tasks requiring global context.
The Bi-Mamba architecture allows us to capture bidirectional contextual relationships by processing sequences in both forward and reverse directions. With the gating mechanism, we can seamlessly integrate these contexts, enabling meaningful applications like cell type annotation and gene interaction prediction. By leveraging a shared-weight design and linear scalability, we enhance the strengths of Mamba, creating a versatile and efficient tool for sequence modeling that aligns with our goals.
3.3 Pretraining Objective
Next Gene Prediction Loss The sequence modeling module processes gene expression sequences to predict the probability distribution of the next gene token conditioned on all previous gene tokens within a cell. The loss function, referred to as $\mathcal{L}_{\text{lang}}$, is computed as the negative log-likelihood (NLL) of the true next gene token $g_{j}$ given the preceding tokens $\{g_{1},g_{2},\dots,g_{j-1}\}$, represented as:
$\mathcal{L}_{\text{lang}}=-\frac{1}{M}\sum_{j=1}^{M}\log P(g_{j}\mid g_{1},g_{2},\dots,g_{j-1})$
This loss function enables end-to-end training of the model by leveraging the sequential representation of gene expression data to predict each gene based on its preceding context within the same cell. By stacking multiple BiMamba layers, the model is trained to capture complex bidirectional relationships among genes, enabling robust analysis and modeling of single-cell gene expression data.
Gene Pathway Loss
To capture the gene-gene relationships within biological pathways, we employ an InfoNCE loss function that enforces similarity among genes sharing a common pathway (Oord et al., 2018; Sharma & Xu, 2023; Gundogdu et al., 2022). Pathways represent functional groupings of genes that collectively contribute to specific biological processes (Zien et al., 2000; García-Campos et al., 2015). Genes within the same pathway are often functionally related and co-regulated, making their representations in the embedding space naturally similar. Conversely, genes that belong to different pathways should be distinct in their embeddings. To achieve this, we define gene pairs within the same pathway as positive pairs (label $y_{ij}=1$) and gene pairs from different pathways as negative pairs (label $y_{ij}=0$).
Given a gene pair $(i,j)$, we compute the cosine similarity between their embeddings:
$\text{sim}(i,j)=\frac{\mathbf{h}_{i}\cdot\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}$
where $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are the normalized embeddings of genes $i$ and $j$. Positive pairs correspond to genes sharing a pathway (label $y_{ij}=1$), while all other gene pairs in the batch are treated as negatives (label $y_{ij}=0$). To emphasize the contrast between positive and negative pairs, we normalize the similarities using temperature scaling:
$\tilde{\text{sim}}(i,j)=\frac{\text{sim}(i,j)}{\tau}$
where $\tau>0$ is a temperature hyperparameter. The InfoNCE loss for a batch of gene pairs is then formulated as:
$\mathcal{L}_{\text{pathway}}=-\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\tilde{\text{sim}}(i,j))}{\sum_{k\in\mathcal{N}(i)}\exp(\tilde{\text{sim}}(i,k))}$
Here, $\mathcal{P}$ denotes the set of positive gene pairs (genes in the same pathway), and $\mathcal{N}(i)$ includes all gene pairs involving gene $i$ in the batch, including negatives. The numerator in the logarithm encourages high similarity for positive pairs, while the denominator includes all other pairs, ensuring the model contrasts between intra-pathway and inter-pathway relationships. This loss function ensures that embeddings of genes within the same pathway are pulled closer together, capturing their shared biological context, while embeddings of genes from different pathways are pushed apart. By leveraging pathway information, this approach enhances the biological interpretability and functional organization of gene embeddings. The final pretraning loss is the weighted sum of the three loss:
$\mathcal{L}=\mathcal{L}_{\text{lang}}+\gamma\mathcal{L}_{\text{pathway}}$
We set the $\gamma$ value to 0.1 in our experiments based on validation results from sample datasets.
4 Experiments
Pretraining Dataset Construction
Our pretrianing dataset was constructed using single-cell RNA sequencing (scRNA-seq) data sourced from the CELLXGENE database, including raw count matrices and their corresponding metadata. The original dataset comprised 50,689,395 cells. To ensure data quality, duplicates were removed by retaining only entries where the feature is_primary_data was set to True. This step eliminated approximately 41% of the samples, resulting in 30,139,066 unique cells. Subsequently, the data was filtered to include 22,053 human protein-coding or miRNA genes. Cells expressing fewer than 200 genes were excluded to remove low-quality samples. The total gene expression counts for each cell were normalized to a fixed target value, and a logarithmic transformation ($log1p$) was applied to scale the data and reduce skewness. Following preprocessing, the dataset used for pretraining was finalized with 29,849,897 cells. For further refinement, a normalization factor was computed for each gene by calculating the non-zero median expression value of that gene across all cells. Genes were then ranked based on their normalized expression levels within each cell, and the top 2,048 or 4,096 gene indices were selected as input features for the pretraining stage.
Pretraining Configuration We configured GeneMamba with 24 Bi-Mamba layers, an inner dimension of 512, and a vocabulary size of 25,426, resulting in 65.74 million learnable parameters. The model was trained for five epochs, distributed across four NVIDIA A100-SXM4-80GB GPUs, requiring approximately three weeks to complete. Additionally, we implemented variations of GeneMamba and compared their performance.
Downstream Task Datasets For the downstream tasks in the GeneMamba paper, we utilized several datasets tailored to specific evaluations. For cell type annotation, we employed hPancreas, MS, Myeloid, and Myeloid_b, with Myeloid_b derived from Myeloid by excluding extremely small cell types to ensure robust performance assessment. Multi-batch analysis was conducted using the PBMC12k, COVID-19, and Perirhinal Cortex datasets to evaluate the model’s effectiveness across diverse batch settings. For gene correlation analysis, the Immune and BMMC datasets were utilized to explore gene-gene relationships and validate the model’s ability to capture meaningful biological insights. More details can be found in Appendix B.1.
Baselines We evaluate our model against established baselines, including GeneFormer, scGPT, scFoundation, and scBert, which are transformer-based models designed for single-cell data analysis. Additionally, we compare with Harmony, a widely used computational biology toolkit for clustering and cell-type classification. These methods provide a comprehensive benchmark, spanning deep learning and traditional computational biology approaches. More details about the baseline models are introduced in Appendix B
4.1 Multi-batch Integration Task
Figure 3: Results of multi-batch integration. Benchmark of the fine-tuned GeneMamba on the PBMC 12k dataset for the multi-batch integration task. The UMAP plot of learned cell embeddings is colored by cell types.
In this experiment, we evaluated the multi-batch integration performance of the GeneMamba model. Multi-batch integration refers to aligning and harmonizing single-cell datasets from multiple experimental batches to minimize batch effects while preserving biologically meaningful patterns. We fine-tuned the pretrained GeneMamba model by adding a simple classification head (MLP) on the PBMC12k, COVID-19, and perirhinal cortex datasets. Using the learned embeddings from these foundation models, we performed multi-batch integration experiments.
The Avg-batch metric assesses the model’s ability to correct batch effects and integrate data across multiple batches, while the Avg-bio metric evaluates the preservation of biological variation and meaningful clustering of cell types after integration. As shown in Table 1 and Figure 3, batch correction and biological preservation often present a trade-off. Harmony, as a specialized method for batch effect correction, excels in eliminating batch effects but tends to ignore biological differences. In contrast, GeneMamba demonstrates superior performance by effectively mitigating batch effects while preserving biological information. This highlights its robust capability in multi-batch integration and its biological relevance for single-cell data analysis. For further details, please refer to Figures 9 and 10.
Table 1: Benchmark results of the models on multi-batch experiments with BATCH and Cell metrics
Metric Dataset Harmony GeneFormer scGPT scFoundation GeneMamba
Avg_batch Immune 0.9514 0.8153 0.9194 0.8904 0.9536
PBMC12k 0.9341 0.9545 0.9755 0.9628 0.9604
BMMC 0.8999 0.7720 0.8431 0.7598 0.9157
Perirhinal Cortex 0.9442 0.9127 0.9600 0.9560 0.9673
Covid-19 0.8781 0.8240 0.8625 0.8346 0.8742
Avg_bio Immune 0.6945 0.6983 0.7879 0.7337 0.8131
PBMC12k 0.7990 0.7891 0.9018 0.8662 0.8344
BMMC 0.6316 0.6324 0.6576 0.5250 0.7628
Perirhinal Cortex 0.8595 0.8547 0.9552 0.9606 0.9062
Covid-19 0.4468 0.5567 0.6476 0.5468 0.5537
4.2 Cell Type Annotation
In this experiment, we evaluate the classification ability of the GeneMamba model on cell type annotation tasks using four benchmark datasets: hPancreas, MS, Myeloid, and Myeloid_b. The Myeloid_b dataset is a modified version of the Myeloid dataset, created by excluding extremely small cell types to ensure a more balanced distribution of cell populations. These datasets encompass diverse biological contexts, allowing for a comprehensive evaluation of the model’s capabilities. Detailed statistics and characteristics of these datasets are provided in Appendix B.1 for reference.
The annotation experiment employs the GeneMamba model fine-tuned with a simple classification head (MLP) in a supervised learning setup, directly testing its ability to learn and predict cell types based on input data. As summarized in Table 2, the GeneMamba model demonstrates competitive performance across datasets. It achieves the highest accuracy (0.9713) and Macro-F1 (0.7710) scores for the hPancreas dataset, showing its robustness in capturing complex cell-type variations. Similarly, in the Myeloid dataset, GeneMamba outperforms others with a Macro-F1 score of 0.3650, showcasing its edge in identifying nuanced differences between cell types. While it lags in the MS dataset, its performance on the balanced Myeloid_b dataset (Acc: 0.9603, Macro-F1: 0.9235) remains strong, underscoring its capability in handling challenging tasks.
These results highlight the effectiveness of GeneMamba’s architecture and training strategy in capturing meaningful biological patterns for precise cell type classification. The embeddings produced by the model capture intrinsic differences across cell types, simplifying downstream tasks and enabling the use of a simple classifier to achieve high accuracy. Furthermore, the confusion matrix in Appendix Figure 11 reveals clear patterns of prediction accuracy and misclassification, reflecting GeneMamba’s robustness and precision in distinguishing between diverse cell types.
Table 2: Benchmark annotation performance across datasets
Datasets Models Acc Macro-F1
hPancreas GeneFormer 0.9665 0.7450
scGPT 0.9710 0.7632
scFoundation 0.9602 0.7101
GeneMamba 0.9713 0.7710
MS GeneFormer 0.7650 0.6220
scGPT 0.8471 0.6630
scFoundation 0.7763 0.6812
GeneMamba 0.6825 0.5342
Myeloid GeneFormer 0.6445 0.3600
scGPT 0.6341 0.3562
scFoundation 0.6446 0.3646
GeneMamba 0.6607 0.3650
Myeloid_b GeneFormer 0.9540 0.9380
scGPT 0.9421 0.9434
scFoundation 0.9574 0.9569
GeneMamba 0.9603 0.9235
4.3 Gene Rank Reconstruction
Figure 4: Gene rank reconstruct results on PBMC12k dataset (a) Venn diagrams showing overlapping between input and output tokens in the pancreas dataset by three models: GeneMamba_U (unidirectional Mamba module as backbone), GeneFormer, GeneMamba (BiMamba module as backbone). (b) Density plots showing input and output ranking in the pancreas dataset by three models: GeneMamba_U, GeneFormer, GeneMamba.
The GeneMamba model demonstrates exceptional capability in reconstructing ranked gene orders, a critical requirement for single-cell analysis tasks where gene expression patterns reveal cellular states and transitions. The analysis begins by randomly selecting sample cells. From these cells, input gene tokens are extracted, and output tokens are generated using the GeneMamba model. Then the gene tokens are ranked based on their predicted likelihoods, with missing tokens filled by zeros to ensure sequence consistency. This approach preserves the structural integrity of the data and facilitates direct comparison between input and output rankings.
Table 3: Gene rank reconstruction performance on the PBMC12k dataset
Model L-Dist BLEU Spearman
GeneMamba_U 430 0.532 0.469
GeneFormer 23 0.968 0.703
GeneMamba 6 0.987 0.711
The consistency of GeneMamba’s predictions is validated through a Venn diagram Figure 4a, which illustrates a high degree of overlap between input and output tokens for the pancreas dataset. This overlap highlights the model’s ability to retain critical features of the input data, an essential characteristic for single-cell studies where the integrity of gene expression ranks directly influences downstream analyses. A density plot Figure 4b provides further evidence of rank fidelity. It demonstrates that lower-ranked input genes consistently yield lower-ranked outputs, indicating a linear relationship between input and output ranks. This observation underscores the model’s effectiveness in capturing patterns within the data and maintaining rank consistency.
Comparative analysis (Figure 4 from left to right) reveals the advantages of GeneMamba over other models: (1) The inclusion of the BiMamba module significantly enhances performance compared to the unidirectional module. This result demonstrates the necessity of bidirectional context extraction for effectively reconstructing ranked gene orders. (2) The GeneFormer model excels at predicting higher-ranked genes but struggles with lower-ranked genes, often treating them as noise due to their lower frequency. In contrast, GeneMamba captures both higher- and lower-ranked genes effectively, showcasing its robustness and sensitivity to varying gene expression levels. The metric results are summarized in Table 3. Additional results are provided in the Appendix Table 7. By reliably reconstructing ranked gene orders, GeneMamba offers a powerful solution for understanding complex gene expression patterns. Its ability to maintain rank integrity across varying data scales positions it as an invaluable tool for single-cell genomics.
4.4 Gene Correlation Analysis
Figure 5: Gene Correlation Analysis. (a) Distribution of consine similarity score for positive and negative gene pairs using Gene2Vec embeddings. (b) Distribution of pearson correlation score for positive and negative gene pairs using Gene2Vec embeddings. (c) Gene-gene topology comparison using adjacency matrices highlights shared and unique structures captured by GeneMamba, scGPT, and Gene2Vec, such as genes RALA, DECR1, and CBL1.
GeneMamba exhibits exceptional capabilities in gene correlation analysis, distinguishing itself in capturing biologically meaningful patterns, maintaining raw data fidelity, and identifying unique model-specific topologies. A series of experiments evaluated GeneMamba’s performance in distinguishing gene relationships, preserving correlation structures, and classifying gene types.
Differentiating Positive and Negative Gene Pairs In the first experiment, GeneMamba’s ability to differentiate between ”positive” and ”negative” gene pairs was assessed using Gene2Vec embeddings. Nearly 30,000 gene pairs, evenly distributed as positive (1) or negative (0), were analyzed for similarity scores using cosine similarity and Pearson correlation. GeneMamba demonstrated a significant separation between the mean similarity scores of positive and negative pairs (Figure 5 a, b). Compared to baseline models, GeneMamba achieved a clearer distinction, emphasizing its capacity to accurately represent gene relationships.
Embedding Alignment To quantify discrimination power, we compared GeneMamba’s embeddings with those of scGPT and scFoundation. The ”distance” between positive and negative pair distributions was measured using Euclidean distance, KL divergence, and JS divergence. The result is shown in Appendix Figure 12. GeneMamba consistently outperformed the other models across all metrics, highlighting its superior ability to distinguish between gene pair types and reinforcing its robustness in modeling gene relationships.
Gene-Gene Topology Analysis The ability to capture gene-gene topologies was analyzed using adjacency matrices derived from embeddings. For the PBMC12K dataset, distances between embeddings of randomly selected gene pairs (100 cells in our case) were calculated and binarized based on threshold value. Then we build the topology graph based on the adjacent matrix (Figure  5c) for each evaluated model. Notably, while identifying unique model-specific topologies, all three models captured biologically significant genes and their interactions, such as RALA, DECR1, and CBL1. This suggests that the foundation models share a similar representation space.
5 Conclusion
In this study, we present GeneMamba, a foundational model designed to advance single-cell analysis. By treating cells as sentences and genes as tokens, and with the BiMamba block and biological aware loss function, GeneMamba achieves a deeper understanding of cellular heterogeneity and gene relationships. This work underscores the potential of adapting large language models for biological data analysis, bridging computational innovation with biological insights. Future directions include enhancing the model with mechanisms like cross-attention for improved transfer learning, scaling to even larger datasets, and creating accessible tools for the research community.
Code Availability
Datasets and code are available at the github address: https://github.com/MineSelf2016/GeneMamba
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. Our work has potential societal implications, but none warrant specific emphasis here.
Appendix
Appendix A Dataset Construction
We constructed our pretraining dataset manually, which is sourced from the cellxgene database, acquiring single-cell RNA sequencing (scRNA-seq) data in raw count matrix format alongside the corresponding metadata. The metadata available on this platform were submitted by the original data contributors, and therefore, we consider it to be partially reliable. It is worth mentioning that, for the first time, we corrected the “organ“ column by mapping it to the “tissue“ column in the original collected data.
The raw data consisted of 50,689,395 cells downloaded from cellxgene. Following the guidelines provided by the platform, which notes that “in some cases, data from the same cell exists in different datasets, therefore cells can be duplicated throughout CELLxGENE Discover and by extension the Census,” we removed duplicates by filtering for entries where the feature is_primary_data was set to True. This step eliminated 20,550,329 samples, leaving a total of 30,139,066 unique cells.
Subsequently, we filtered the data to include only 22,053 human protein-coding or miRNA genes. To ensure data quality and reduce noise from poor-quality cells, we retained cells expressing at least 200 genes. The total counts of gene expression values in each cell were then normalized to a fixed target value to ensure comparability across cells. Following normalization, we applied a logarithmic transformation (log1p) to scale the data for the following analyses.
After preprocessing, the dataset for pretraining was finalized with 29,849,897 cells. We calculated a normalization factor for each gene by determining the non-zero median expression value of that gene across all cells. This normalization factor was consistently applied to both the pretraining corpus and all future datasets presented to the model. This approach deprioritized ubiquitously highly expressed housekeeping genes while highlighting genes with lower expression levels that contribute significantly to distinguishing cell states.
For the tokenization process, genes were ranked by their normalized expression values within each cell, and the resulting tokenized data were stored in the Huggingface Datasets format, retaining all non-zero gene entries. To provide a non-parametric representation of the transcriptome for each cell, we selected the top 2,048 or 4,096 gene indices as the input data to pretraining stage.
Appendix B Downstream Datasets, Tasks and Evaluation Metrics
B.1 Downstream Datasets
The statistics of the datasets are shown in the Table  4.
Table 4: Statistics of downstream datasets, including the number of cells (n_cells), genes (n_genes), batches (n_batches), and cell types (n_cell_types) for each dataset.
Dataset n_cells n_genes n_batches n_cell_types
bmmc 90261 14087 12 45
covid19 20000 1200 2 39
pbmc12k 11990 3346 2 9
perirhinal cortex 17535 59357 2 10
immune 33506 12303 10 16
myeloid 13178 3000 2 21
hpancreas 14818 3000 2 14
ms 21312 3000 2 18
hPancreas This dataset comprises single-cell RNA sequencing (scRNA-seq) data from human pancreatic cells, including various cell types such as alpha, beta, delta, and acinar cells. It is utilized to study cellular heterogeneity and gene expression profiles within the pancreas.
MS The multiple sclerosis (MS) dataset includes scRNA-seq data from peripheral blood mononuclear cells (PBMCs) of individuals with MS. It aids in understanding immune cell composition and gene expression changes associated with MS.
Myeloid This dataset contains scRNA-seq data focusing on myeloid cells, such as monocytes and macrophages, from various tissues or conditions. It is essential for studying the role of myeloid cells in immune responses and diseases.
PBMC12k The PBMC12k dataset consists of scRNA-seq data from 12,000 peripheral blood mononuclear cells obtained from a healthy donor. It serves as a reference for immune cell types and their gene expression profiles in the bloodstream.
BMMC This dataset includes scRNA-seq data from bone marrow mononuclear cells, encompassing various hematopoietic cell types. It is utilized to study hematopoiesis and bone marrow microenvironments.
COVID19 The COVID19 dataset comprises scRNA-seq data from PBMCs of COVID-19 patients. It facilitates the investigation of immune responses and cellular changes during SARS-CoV-2 infection.
Immune Human This dataset contains scRNA-seq data from human immune cells across different tissues or conditions. It aids in understanding the diversity and function of the human immune system.
Perirhinal Cortex This dataset includes scRNA-seq data from cells of the perirhinal cortex, a region of the brain involved in memory and recognition. It is used to study neuronal and glial cell types and their gene expression profiles in this specific brain area.
The original Myeloid dataset is highly imbalanced. To better evaluate the models’ performance, we excluded rare cell types (proportion $<$ 0.05) to create the Myeloid_b dataset. Figure 6 illustrates the cell type distribution in both the original and modified Myeloid datasets.
Figure 6: Cell type distribution in the original and modified Myeloid datasets
B.2 Downstream Tasks
We standardize the downstream datasets by converting the batch key to ”batch” and the label key to ”celltype.” We then use a manual stratified split to divide the dataset into train/test with a 0.9/0.1 ratio to avoid missing rare classes in the test set. For example, in the COVID dataset, using ‘train_test_split()‘ directly would cause the 22nd class to be missing, as it contains only 3 samples, and a 0.3 test sample rate would result in zero samples, causing errors during AUC-ROC calculation. The split column is labeled ”partition,” where ”train” is for training and ”test” is for testing.
Multi-batch integration Task
Multi-batch integration is a task in single-cell data analysis that involves harmonizing data collected from multiple experimental batches to mitigate technical differences while preserving meaningful biological information. In this context, let $X=\{X_{1},X_{2},\ldots,X_{n}\}$ denote datasets from $n$ distinct batches, where $X_{i}\in\mathbb{R}^{m_{i}\times p}$ represents the gene expression matrix of the $i$-th batch with $m_{i}$ cells and $p$ genes. The objective is to project these datasets into a unified latent space $Z\in\mathbb{R}^{M\times d}$, where $M=\sum_{i=1}^{n}m_{i}$ is the total number of cells across all batches, and $d$ is the dimensionality of the latent space. In this shared space, cells with similar biological profiles are grouped together, regardless of their batch origin. This task is crucial for ensuring the comparability of cells from different batches and facilitating downstream analyses like clustering or cell type classification. Successful integration minimizes batch-specific artifacts and emphasizes biologically relevant signals, enabling a coherent and unbiased interpretation of single-cell data.
Cell Type Annotation
Cell type annotation is a fundamental task in single-cell transcriptomics aimed at assigning biologically meaningful labels to individual cells based on their gene expression profiles. Formally, given a gene expression matrix $X\in\mathbb{R}^{M\times p}$, where $M$ is the total number of cells and $p$ represents the number of genes, the goal is to map each cell $x_{i}\in\mathbb{R}^{p}$ to a cell type label $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{k}\}$ is the set of $k$ predefined cell types. This task leverages the distinct transcriptional signatures of cell types, often characterized by differential expression of marker genes, to provide insights into cellular identity and function. Successful annotation ensures that biologically relevant features are preserved, enabling accurate downstream analyses such as the study of cellular heterogeneity, lineage tracing, and disease-specific cell state identification.
Gene Rank Reconstruction
Figure 7: Gene Rank Reconstruct framework of GeneMamba Gene Imputation framework evaluating rank reconstruction by predicting (n+1)-th tokens iteratively.
In the figure 7, the framework of doing the gene rank reconstruction experiment. The framework demonstrates the iterative nature of rank reconstruction, where the model processes one token at a time, maintaining sequential dependencies inherent in biological systems. This approach mirrors natural biological progressions, further emphasizing the model’s relevance to genomics.
Gene Distribution Alignment
In this experiment, we evaluate the GeneMamba’s performance to distinguish between positive pairs and negative pairs, the reference data is collected in the Gene2Vec paper (Du et al., 2019). And we generate the embeddings using different models for positive gene pairs and negative gene pairs, then we measure the embeddings similarity of each pair. We evaluate how the distribution of positive pairs and negative pairs differ by calculating the eucilean distance, KL-Divergence, and JS-Divergence of the two distributions, the result is shown in figure 5.
Gene Topology Preservation
To evaluate gene-gene topology preservation, we conducted experiments using the pbmc12k dataset. A random selection of gene-gene pairs was made from the dataset, and their embeddings were obtained using three models: scGPT, GeneMamba, and Gene2Vec. The Euclidean distances between the gene pairs were computed, and an adjacency matrix was constructed to represent these relationships. The threshold for the distances was set as the mean value of the matrix elements, and the distances were binarized into 0 and 1 based on this threshold. To compare the topological similarity among models, the Jaccard distances between the adjacency matrices were calculated. Additionally, overlapping gene-gene pairs were identified to further understand shared topology. The experiment culminated in a visual representation of the gene-gene topology, highlighting the models’ ability to preserve biological relationships in gene expression data.
B.3 Evaluation Metrics
AvgBIO
To evaluate how well the integration preserves biological signals, we calculate the AvgBIO metric. This metric is the average of three biological conservation measures: the Adjusted Rand Index ($ARI_{\text{cell}}$), Normalized Mutual Information ($NMI_{\text{cell}}$), and the Average Silhouette Width based on cell types ($ASW_{\text{cell}}$):
$\text{AvgBIO}=\frac{ARI_{\text{cell}}+NMI_{\text{cell}}+ASW_{\text{cell}}}{3}.$
1. Adjusted Rand Index ($ARI_{\text{cell}}$): We use $ARI_{\text{cell}}$ to quantify the agreement between the true biological labels and the clusters predicted after integration. By adjusting for chance agreement, $ARI_{\text{cell}}$ captures how well the integration maintains the clustering structure:
$ARI_{\text{cell}}=\frac{\text{Index observed}-\text{Index expected}}{\text{Maxindex}-\text{Index expected}}.$
Here, the observed index measures the agreement between clustering results and the ground truth, while the expected index accounts for random chance. We interpret $ARI_{\text{cell}}$ scores ranging from 0 (random labeling) to 1 (perfect agreement).
2. Normalized Mutual Information ($NMI_{\text{cell}}$): To evaluate how much information is shared between the true biological labels ($Y$) and the predicted cluster labels ($C$), we compute $NMI_{\text{cell}}$ using:
$NMI_{\text{cell}}=\frac{2\cdot I(Y;C)}{H(Y)+H(C)},$
where $I(Y;C)$ is the mutual information, and $H(Y)$ and $H(C)$ are the entropies of the true and predicted labels. This score ranges from 0 (no alignment) to 1 (perfect alignment), and we use it to assess the consistency of clustering.
3. Average Silhouette Width ($ASW_{\text{cell}}$): We calculate $ASW_{\text{cell}}$ to measure how well-separated clusters are based on cell type labels. The silhouette score ($ASW_{C}$) evaluates whether cells are closer to their own cluster than to other clusters. We normalize the silhouette score using:
$ASW_{\text{cell}}=\frac{ASW_{C}+1}{2}.$
This normalization ensures that the score ranges between 0 (poor separation) and 1 (perfect separation).
AvgBatch
To assess how well the integration removes batch effects, we calculate AvgBatch as the average of two metrics: the Average Silhouette Width for batch labels ($ASW_{\text{batch}}$) and Graph Connectivity ($GraphConn$):
$\text{AvgBatch}=\frac{ASW_{\text{batch}}+GraphConn}{2}.$
1. Average Silhouette Width for Batch Labels ($ASW_{\text{batch}}$): To evaluate the extent of batch effect removal, we compute $ASW_{\text{batch}}$. First, we calculate the silhouette score based on batch labels ($ASW_{B}$), which measures how batch-specific artifacts affect the integrated space. Then, we adjust it to reflect batch mixing:
$ASW_{\text{batch}}=1-|ASW_{B}|.$
A higher $ASW_{\text{batch}}$ score indicates better mixing of batches in the latent space, as cells are distributed independently of their batch origin.
2. Graph Connectivity ($GraphConn$): To measure the connectivity of cells within the same biological type, we construct a k-nearest neighbors (kNN) graph for each cell type. Then, we identify the largest connected component ($LCC$) within each graph and calculate the connectivity score as:
$GraphConn=\frac{1}{|C|}\sum_{c\in C}\frac{|LCC(G_{c}^{\text{kNN}})|}{N_{c}}.$
Here, $C$ is the set of cell types, $|LCC(G_{c}^{\text{kNN}})|$ is the size of the largest connected component for cell type $c$, and $N_{c}$ is the total number of cells in $c$. We use this score to evaluate whether cells of the same type remain well-connected after integration.
In the reconstruct experiments, we use Exact Match, Levenshtein Distance, and BLEU Score to evaluate the reconstruction ability of the GeneMamba model. We refer to the input gene token sequence as $S_{\text{in}}$ and the model output sequence as $S_{\text{out}}$. These metrics are defined as follows:
Exact Match
The Exact Match (EM) measures whether the input sequence $S_{\text{in}}$ is identical to the output sequence $S_{\text{out}}$. It is defined as:
$\text{EM}=\begin{cases}1,&\text{if }S_{\text{in}}=S_{\text{out}},\\ 0,&\text{otherwise}.\end{cases}$
For a dataset with $N$ cells, the overall Exact Match score is the average:
$\text{EM}_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}\text{EM}(S_{\text{in}}^{i},S_{\text{out}}^{i}).$
Levenshtein Distance
The Levenshtein Distance (LD) quantifies the minimum number of single-character edits (insertions, deletions, or substitutions) required to transform $S_{\text{in}}$ into $S_{\text{out}}$. Formally:
$\text{LD}(S_{\text{in}},S_{\text{out}})=\min\begin{cases}\text{LD}(S_{\text{in}}[:i-1],S_{\text{out}}[:j])+1,&\text{deletion},\\ \text{LD}(S_{\text{in}}[:i],S_{\text{out}}[:j-1])+1,&\text{insertion},\\ \text{LD}(S_{\text{in}}[:i-1],S_{\text{out}}[:j-1])+c,&\text{substitution},\end{cases}$
where $c=0$ if $S_{\text{in}}[i]=S_{\text{out}}[j]$, otherwise $c=1$.
For normalization across sequences of varying lengths, the Normalized Levenshtein Distance is computed as:
$\text{NLD}=1-\frac{\text{LD}(S_{\text{in}},S_{\text{out}})}{\max(|S_{\text{in}}|,|S_{\text{out}}|)},$
where $|S|$ denotes the length of sequence $S$.
BLEU Score
The BLEU Score (Bilingual Evaluation Understudy) evaluates the similarity between $S_{\text{out}}$ and $S_{\text{in}}$ by comparing n-grams. It is computed as:
$\text{BLEU}=\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)\cdot\text{BP},$
where:
  • $p_{n}$ is the precision of $n$-grams,
  • $w_{n}$ is the weight for $n$-grams (often $w_{n}=\frac{1}{N}$ for uniform weights),
  • BP is the brevity penalty to penalize short outputs.
The brevity penalty is defined as:
$\text{BP}=\begin{cases}1,&\text{if }|S_{\text{out}}|\geq|S_{\text{in}}|,\\ \exp(1-\frac{|S_{\text{in}}|}{|S_{\text{out}}|}),&\text{otherwise}.\end{cases}$
The BLEU score ranges from 0 (no similarity) to 1 (perfect match).
These metrics together provide a comprehensive assessment of the GeneMamba model’s ability to reconstruct input gene token sequences accurately.
Baselines
To evaluate the performance of our proposed approach, we compare it with several state-of-the-art baseline methods. Below, we briefly describe each method:
  • GeneFormer (Theodoris et al., 2023): A transformer-based model specifically designed for single-cell analysis. It uses a gene-token representation to capture complex gene-gene relationships and downstream cellular heterogeneity.
  • scGPT (Cui et al., 2024a): A generative pre-trained transformer adapted for single-cell data analysis. scGPT utilizes a pretraining-finetuning paradigm to handle various single-cell tasks, including cell-type classification and clustering.
  • scFoundation (Hao et al., 2024): A foundational model for single-cell data analysis that incorporates self-supervised learning techniques to generate robust embeddings for single-cell profiles. It excels in tasks requiring integration across batches and datasets.
  • scBERT (Yang et al., 2022): A BERT-like architecture adapted for single-cell data. scBERT leverages bidirectional contextual embeddings to model complex interactions between genes within cells, making it suitable for both classification and clustering tasks.
  • Harmony (Korsunsky et al., 2019): A batch-effect correction method that aligns single-cell datasets from different conditions or experiments. This method is often used as a preprocessing step to ensure integration and compatibility across datasets before applying downstream machine learning methods.
Finetuning Details
For the cell type annotation task, we append the [CLS] token and add a multi-layer perceptron (MLP) to further train the model on the downstream task in a supervised manner. The train/test split is predefined by the dataset provider, so we do not perform any additional splitting. We use the generated [CLS] token embedding as the representation of the cell.
For the multi-batch integration task, we utilize the model fine-tuned on the cell type annotation task, remove the MLP layer, and directly extract embeddings from the GeneMamba model to conduct the experiment.
Gene Analysis Experiment
The gene embeddings for different models are obtained as follows:
  • Gene2Vec: The embeddings are generated using the Gene2Vec model, where each gene corresponds to a unique embedding.
  • scGPT: The embeddings are extracted using the provided function of scGPT, with each gene mapped to a distinct embedding.
  • scFoundation: Since the scFoundation model does not provide a direct method for obtaining gene embeddings, we apply mean pooling to the generated context-aware embeddings of scFoundation on downstream datasets.
Appendix C More Related Work
C.1 Discretization of Gene Expression
The transformation of gene expression levels into machine-readable tokens remains a critical aspect of single-cell modeling, with emerging methods aiming to balance biological relevance and computational efficiency.
BinBased Strategy: Binning is the most commonly employed strategy for discretization, where continuous gene expression values are converted into categorical tokens, allowing transformers to process the data. Two main binning strategies are widely used:
Fixed-Size Binning: As seen in scBERT, fixed-size binning divides the range of expression values into equally sized intervals. Genes falling within the same interval are represented identically. This method, however, often leads to significant information loss, as most genes with low expression levels cluster within the same bin (e.g., the [0,1) bin (Boiarsky et al., 2023)), masking subtle but biologically relevant differences.
Adaptive Binning: Adaptive binning, employed by scGPT, dynamically adjusts bin sizes such that each interval represents an equal proportion of expressed genes within a single cell. While this approach aims to better capture the distribution of gene expression, it still faces challenges related to information loss, particularly for genes with extreme expression values or biological significance.
Value Projection Strategy: The value projection strategy, first introduced by TOSICA (Chen et al., 2023), has since been adopted by scCLIP (Xiong et al., 2023), SpaFormer (Wen et al., 2023b), CellLM (Zhao et al., 2023), and CellPLM (Wen et al., 2023a). In this approach, continuous gene expression values are directly projected into high-dimensional embeddings through linear or non-linear transformations. This method retains the full resolution of the data, avoiding the information loss associated with binning. However, feeding continuous values into transformer-based models diverges from traditional tokenization approaches in NLP, potentially impacting model performance. Additionally, continuous embeddings may be more sensitive to noise and batch effects, necessitating robust preprocessing steps.
Rank-Based Strategy: The rank-based strategy was first introduced in iSEEEK (Shen et al., 2022) for single-cell transformers. This approach ranks gene expression values within each cell and retains only the top-ranked genes. The remaining genes are either truncated or represented as a lower-priority group. By focusing on the top-ranked genes, this strategy reduces sequence length, significantly improving computational efficiency. It also aligns well with the biological prioritization of highly expressed genes in specific cell states or conditions.
C.2 Model Architectures for Single-Cell Analysis
Recent advancements in model architectures have addressed the computational challenges of transformer-based models in single-cell RNA sequencing (scRNA-seq) analysis, with notable innovations including scTCA (Yu et al., 2024), a hybrid Transformer-CNN architecture designed for imputation and denoising of single-cell read counts to enhance downstream analyses; scHyena (Oh et al., 2023), a foundation model for full-length scRNA-seq analysis in brain tissues that uses a novel transformer architecture to process data without information loss, improving tasks like cell type classification and data imputation; xTrimoGene (Gong et al., 2024b), an efficient and scalable representation learner employing an asymmetric encoder-decoder transformer architecture to reduce computational requirements while maintaining high accuracy in tasks such as cell type annotation and drug combination prediction; scTransSort (Jiao et al., 2023), a transformer-based model for intelligent cell type annotation that leverages self-attention mechanisms to extract features from scRNA-seq data, achieving high accuracy and performance; scGraphformer (Fan et al., 2024), which employs transformer-based graph neural networks to provide accurate and scalable cell type annotations, uncovering cellular heterogeneity and interactions in scRNA-seq data; scGFT (Nouri, 2025), a train-free, cell-centric generative model adept at synthesizing scRNA-seq data to enhance data augmentation and analysis; White-Box Diffusion Transformer(Cui et al., 2024b), a hybrid model combining diffusion models and white-box transformers to generate synthetic and biologically plausible scRNA-seq data, improving training efficiency and resource utilization; STGRNS (Xu et al., 2023), an interpretable transformer-based method for inferring gene regulatory networks from scRNA-seq data to enhance understanding of gene interactions; scRDiT (Dong et al., 2024), a generative approach using diffusion transformers to create virtual scRNA-seq data, facilitating the study of gene expression dynamics; and TransformerST (Zhao et al., 2024a), an unsupervised model based on the transformer architecture for super-resolution in spatial transcriptomics, improving spatial gene expression analysis. These models collectively advance the field by addressing key challenges in scRNA-seq data processing and interpretation.
Traditional single-cell large language models (LLMs) are predominantly based on transformer architectures (Boiarsky et al., 2023; White & Harris, 2024). While transformers have demonstrated exceptional performance in various single-cell transcriptomics tasks, their high computational and memory requirements make training large models time-intensive and resource-heavy. These limitations become particularly evident when processing the ultra-long sequences typical of single-cell RNA-seq data. To alleviate these challenges, some transformer-based models have incorporated optimizations to improve efficiency. For example, scBERT integrates Performer (Choromanski et al., 2020) to approximate attention calculations, while scGPT employs Flash Attention (Dao et al., 2022) to accelerate training by reducing memory overhead. Despite these advancements, the overall time required to train such models remains substantial, especially for datasets with millions of cells and thousands of genes. To address the intrinsic limitations of transformers, Bidirectional Mamba (Bi-Mamba) has been proposed as a novel alternative architecture. Bi-Mamba leverages state-space models (SSMs) to process ultra-long sequences with linear computational complexity in both time and memory. This represents a significant departure from traditional transformer-based methods, offering a scalable and efficient solution for the high-dimensional, high-throughput nature of single-cell transcriptomics data.
Appendix D Discussion of Scalability
Scalability is an essential feature for models designed to handle large-scale single-cell omics datasets. The GeneMamba model demonstrates superior scalability and computational efficiency across various configurations, including sequence lengths and model complexities. In this section, we discuss its scalability based on the results summarized in Tables 5 and 6.
Training Configurations and Model Complexity
Table 5 presents the training configurations for the Mamba and GeneMamba models, highlighting the relationship between model complexity (e.g., number of layers, hidden size) and training requirements. For instance, increasing the number of layers from 24 to 48 doubles the parameters and correspondingly increases training time two times (from 5.5 hours to 11 hours). Additionally, GeneMamba demonstrates its capability to handle longer sequence lengths (e.g., 4096 tokens) with a linear increase in training time (from 5.5 hours to 11 hours), showcasing its scalability for more input gene tokens of single-cell data.
Mamba Layers Hidden Size Seq Length Number of Params Training Time (hours/m)
24 512 2048 65.74M 5.5
48 512 2048 105.42M 10.5
24 768 2048 127.05M 11
48 768 2048 215.03M 22
24 512 4096 65.74M 11
Table 5: Training configurations for the GeneMamba models, demonstrating scalability across layers, hidden sizes, and sequence lengths. The data highlights the increasing number of parameters and training time as model complexity grows, with GeneMamba showing efficient handling of longer sequence lengths (e.g., 4096) compared to traditional Mamba setups.
Computational Efficiency Comparison
Table 6 compares the computational efficiency of the Transformer and GeneMamba backbones. GeneMamba requires significantly fewer FLOPs per sample and achieves faster training times across all sequence lengths. For example, at a sequence length of 2048, GeneMamba’s training time is 0.1284 seconds/sample compared to the Transformer’s 0.2227 seconds/sample, a reduction of approximately 42%. This efficiency becomes even more pronounced at longer sequence lengths.
Model Backbone Parameters Forward + Backward FLOPs/sample Seq Length Training Time (seconds/sample)
Transformer 130M 2.85E+12 2048 0.2227
Transformer 130M 7.18E+12 4096 0.6025
Transformer 130M 2.02E+13 8192 1.8904
GeneMamba 130M 1.58E+12 2048 0.1284
GeneMamba 130M 3.15E+12 4096 0.2393
GeneMamba 130M 6.31E+12 8192 0.4885
Table 6: Computational efficiency comparison of Transformer and GeneMamba backbones at different input sequence lengths. GeneMamba significantly reduces computational cost while maintaining high efficiency.
Theoretical Computation and Demonstration
Using the 2048-sequence-length GeneMamba model as an example, the total FLOPs required for processing 2 million cells can be calculated as:
Total FLOPs $\text{Total FLOPs required}=1.58\times 10^{12}\times 2\times 10^{6}=3.16\times 10^{18}\text{ FLOPs.}$.
Given that the A100 80GB GPU has a peak performance of 19.5 TFLOPS for single-precision (FP32) and 312 TFLOPS for half-precision (FP16), the theoretical training time is computed as follows:
For FP32:
$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{19.5\times 10^{12}\text{ FLOPs/sec}}\approx 1.621\times 10^{5}\text{ seconds}=45\text{ hours.}$.
For FP16:
$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{312\times 10^{12}\text{ FLOPs/sec}}\approx 1.013\times 10^{4}\text{ seconds}=2.8\text{ hours.}$.
These results demonstrate the substantial efficiency improvements provided by GeneMamba, especially when leveraging FP16 precision, which reduces training time from 45 hours to just 2.8 hours.
Conclusion on Scalability
GeneMamba’s ability to process longer sequences with lower computational cost and its effective utilization of GPU resources make it a scalable solution for large-scale single-cell analysis. As shown in Tables 5 and 6, its scalability and efficiency significantly outperform traditional Transformer architectures, providing a robust framework for computational biology and related applications.
Appendix E Extended Figures
Figure 8: Results of cell type annotation. Bar plot of the classification metrics across various datasets of GeneMamba.
Figure 9: Results of multi-batch integration. Benchmark of the fine-tuned GeneMamba on the Covid19 dataset for the multi-batch integration task. The UMAP plot of learned cell embeddings is colored by cell types.
Figure 10: Results of multi-batch integration. Benchmark of the fine-tuned GeneMamba on the Perirhinal Cortex dataset for the multi-batch integration task. The UMAP plot of learned cell embeddings is colored by cell types.
Figure 11: Cell type annotation results of GeneMamba. a. Confusion matrix of dataset hPancreas. b. Confusion matrix of dataset MS. c. Confusion matrix of dataset Myeloid. d. Confusion matrix of dataset Myeloid_b.
Figure 12: Gene-gene pairs correlation analysis. The Euclidean Distance, KL Divergence, and JS Divergence across three embedding methods (Gene2Vec, GeneMamba, and Random). A lower value indicates higher similarity. The embeddings generated by GeneMamba are more similar to Gene2Vec embeddings compared to randomly generated embeddings.
Appendix F Extended Tables
Table 7: Gene Rank Reconstruction Performance comparison of GeneMamba, GeneFormer, and GeneMamba_S models across datasets.
Datasets Models L-Dist BLEU Spearman
PBMC12k GeneMamba_U 430 0.532 0.469
GeneFormer 23 0.968 0.703
GeneMamba 6 0.987 0.711
Pancreas GeneMamba_U 370 0.524 0.461
GeneFormer 25 0.956 0.763
GeneMamba 12 0.991 0.792
Zheng68k GeneMamba_U 432 0.581 0.503
GeneFormer 25 0.937 0.901
GeneMamba 11 0.996 0.980
Immune GeneMamba_U 468 0.659 0.442
GeneFormer 17 0.962 0.823
GeneMamba 12 0.998 0.844

### v2

@@
-GeneMamba: Efficient and Effective Foundation Model on Single Cell Data
+GeneMamba: An Efficient and Effective Foundation Model on Single Cell Data
 Abstract
 Transformer-based architectures have shown promise across various domains but struggle with computational inefficiency and scalability.
-To address these challenges, we introduce GeneMamba, a novel model designed specifically for single-cell data analysis.
-GeneMamba incorporates the BiMamba module to efficiently capture gene context information and employs biologically meaningful loss functions during training.
+To address these challenges, we introduce GeneMamba, a novel model specifically designed for single-cell data analysis.
+GeneMamba incorporates the BiMamba module to efficiently capture gene context information and employs biologically informed loss functions during training.
 The model enables scalable processing of over 50 million cells while significantly reducing computational costs.
@@
 Furthermore, reconstruction experiments highlight GeneMamba’s explainability, establishing it as a robust foundation for advancing single-cell transcriptomics in biological and biomedical research.
+By unifying bidirectional state space dynamics with rank-based gene tokenization, GeneMamba achieves both contextual depth and biological interpretability.
+These capabilities position GeneMamba as a scalable and principled backbone for future foundation models in single-cell omics.
 1 Introduction
-Single-cell RNA sequencing (scRNA-seq) has fundamentally transformed our ability to study cellular heterogeneity and dynamics by enabling high-resolution profiling of gene expression at the individual cell level (Korsunsky et al., 2019; Theodoris et al., 2023; Smith & Doe, 2024).
-This technology allows researchers to dissect complex cellular compositions within tissues, trace differentiation trajectories, and identify rare cell populations that remain undetectable with bulk RNA sequencing (Shen et al., 2023; Wen et al., 2023b; Nguyen & Tran, 2024).
-The rich transcriptomic insights provided by scRNA-seq have catalyzed significant advancements in developmental biology, disease modeling, and drug discovery (Gu & Dao, 2023; Yang et al., 2024; Li & Wang, 2024).
-Despite these advantages, the computational analysis of scRNA-seq data presents formidable challenges due to its high dimensionality, inherent sparsity, and technical noise (Du et al., 2019; Zhao et al., 2023; Brown & Davis, 2024).
+Single-cell RNA sequencing (scRNA-seq) has fundamentally transformed our ability to study cellular heterogeneity and dynamics by enabling high-resolution profiling of gene expression at the individual cell level korsunsky2019fast; theodoris2023transfer.
+This technology allows researchers to dissect complex cellular compositions within tissues, trace differentiation trajectories, and identify rare cell populations that remain undetectable with bulk RNA sequencing shen2023generative; wen2023single; deconinck2021recent.
+The rich transcriptomic insights provided by scRNA-seq have catalyzed significant advancements in developmental biology, disease modeling, and drug discovery gu2023mamba; yang2024genecompass; li2024single.
+Despite these advantages, the computational analysis of scRNA-seq data presents formidable challenges due to its high dimensionality, inherent sparsity, and technical noise du2019gene2vec; zhao2023large.
 Addressing these challenges requires robust computational models capable of capturing biologically meaningful patterns while efficiently handling the scale and variability of single-cell data.
-Transformer-based architectures, such as scBERT (Yang et al., 2022) and scGPT (Cui et al., 2024a), have emerged as powerful tools in this domain, demonstrating strong performance in tasks such as cell type classification, gene expression imputation, and differential expression analysis (Bian et al., 2024; Wen et al., 2023a; Garcia & Thompson, 2024).
+Transformer-based architectures, such as scBERT yang2022scbert and scGPT cui2024scgpt, have emerged as powerful tools in this domain, demonstrating strong performance in tasks such as cell type classification, gene expression imputation, and differential expression analysis bian2024scmulan; wen2023cellplm; szalata2024transformers.
 However, transformers exhibit critical limitations when applied to scRNA-seq data.
-The quadratic complexity of their self-attention mechanism constrains their scalability for long sequences, which are characteristic of single-cell transcriptomes (Vaswani, 2017; Choromanski et al., 2020; Smith & Doe, 2024).
-Moreover, transformers often struggle to effectively capture long-range dependencies, which are essential for modeling gene regulatory interactions and cell state transitions (Dao et al., 2022; Chen et al., 2023).
-These limitations have driven the exploration of alternative architectures, among which state space models (SSMs) have emerged as promising solutions, offering improved efficiency and scalability for processing long sequences (Gu & Dao, 2023; Dao & Gu, 2024; Martinez & Hernandez, 2024).
+The quadratic complexity of their self-attention mechanism constrains their scalability for long sequences, which are characteristic of single-cell transcriptomes vaswani2017attention; choromanski2020rethinking.
+Moreover, transformers often struggle to effectively capture long-range dependencies, which are essential for modeling gene regulatory interactions and cell state transitions dao2022flashattention; chen2023transformer.
+These limitations have driven the exploration of alternative architectures, among which state space models (SSMs) have emerged as promising solutions, offering improved efficiency and scalability for processing long sequences gu2023mamba; mamba2.
 In this study, we introduce GeneMamba, a novel state space model designed to efficiently train large-scale cell models on scRNA-seq data.
-SSM-based models have demonstrated competitive or superior performance compared to transformers while significantly reducing computational overhead (Shen et al., 2022; Gong et al., 2024a; Lee & Kim, 2024).
-Building upon this foundation, GeneMamba incorporates bidirectional computation, enhancing its ability to capture both upstream and downstream contextual dependencies, thereby improving its performance in tasks requiring global contextual awareness (Liu et al., 2024; Liang et al., 2024).
+SSM-based models have demonstrated competitive or superior performance compared to transformers while significantly reducing computational overhead shen2022universal; gong2024xtrimogene; ryu2023integration.
+Building upon this foundation, GeneMamba incorporates bidirectional computation, enhancing its ability to capture both upstream and downstream contextual dependencies, thereby improving its performance in tasks requiring global contextual awareness liu2024bidirectional; liang2024bi.
 We validate GeneMamba across a diverse set of applications, including multi-batch integration, cell type annotation, and gene-gene correlation analysis.
-Our experimental results demonstrate substantial improvements in computational efficiency and predictive performance compared to existing models (Hao et al., 2024; Sun et al., 2024; Johnson & Lee, 2024).
-Furthermore, to assess its scalability, we conduct a comprehensive performance analysis on various variations of GeneMamba (Appendix D), highlighting its advantages in real-world biological applications.
+Our experimental results demonstrate substantial improvements in computational efficiency and predictive performance compared to existing models hao2024large; sun2024bidirectional; petegrosso2020machine.
+Furthermore, to assess its scalability, we conduct a comprehensive performance analysis on various variations of GeneMamba (Appendix F), highlighting its advantages in real-world biological applications.
 Our contributions are threefold:
@@
 Existing methods employ various tokenization strategies, each with distinct advantages and limitations.
-Bin-based discretization, as used by scBERT (Yang et al., 2022), scGPT (Cui et al., 2024a), and scMulan (Bian et al., 2024), groups expression values into predefined bins.
+Bin-based discretization, as used by scBERT yang2022scbert, scGPT cui2024scgpt, and scMulan bian2024scmulan, groups expression values into predefined bins.
 This approach preserves absolute value distributions and simplifies sequence modeling, but may introduce information loss, particularly for genes with subtle but biologically significant expression differences.
 Additionally, binning can be sensitive to parameter selection, affecting downstream results.
-Value projection(Szałata et al., 2024), adopted by scFoundation (Hao et al., 2024) and its backbone model xTrimoGene (Gong et al., 2024a), projects gene expression values into continuous embeddings rather than discrete categories.
+Value projectionszalata2024transformers, adopted by scFoundation hao2024large and its backbone model xTrimoGene gong2024xtrimogene, projects gene expression values into continuous embeddings rather than discrete categories.
 This method maintains full data resolution by applying a linear transformation to the gene expression vector, which is then combined with gene-specific embeddings.
 However, the use of continuous embeddings diverges from traditional tokenization strategies in NLP-based transformers, and its impact on model performance remains an open question.
-Rank-based discretization, utilized by Geneformer (Theodoris et al., 2023), GeneCompass (Yang et al., 2024), tGPT(Shen et al., 2023) and LangCell (Zhao et al., 2024b), transforms gene expression values into ordinal rankings.
+Rank-based discretization, utilized by Geneformer theodoris2023transfer, GeneCompass yang2024genecompass, tGPTshen2023generative and LangCell zhao2024langcell, transforms gene expression values into ordinal rankings.
 This approach effectively captures relative expression levels and is more robust to batch effects and noise.
@@
 These models have shown promising results in various tasks, including cell-type annotation, batch integration, and multiomics analysis.
-However, transformer(Vaswani, 2017) has inherent limitations.
+However, transformervaswani2017attention has inherent limitations.
 The most significant one is their quadratic computational complexity with respect to sequence length, which makes them less feasible for long sequences typical of scRNA-seq data.
@@
 Additionally, transformer often struggles with capturing long-range dependencies in sequences, which is crucial for understanding gene regulatory networks and cell state transitions.
-Mamba(Gu & Dao, 2023) has been introduced to address the transformer efficiency problem.
-Bidirectional Mamba(Bi-Mamba) has recently been used to solve long-sequence problems in tasks such as feature extraction(Sun et al., 2024), sequential recommendation(Liu et al., 2024), and time series forecasting(Liang et al., 2024).
+Mambagu2023mamba has been introduced to address the transformer efficiency problem.
+Bidirectional Mamba(Bi-Mamba) has recently been used to solve long-sequence problems in tasks such as feature extractionsun2024bidirectional, sequential recommendationliu2024bidirectional, and time series forecastingliang2024bi.
 Bi-Mamba is designed to efficiently process ultra-long sequences with linear computational complexity, offering a significant reduction in memory and computation requirements compared to transformer.
@@
 This architecture represents a promising alternative to transformer-based methods, addressing their key limitations while maintaining high performance in various single-cell analysis tasks.
-More related work is provied in Appendix C.
+More related work is provided in Appendix E.
 3 Methods
@@
 Once pretrained, the model and its embeddings are employed for various downstream tasks to evaluate the model’s performance.
-Figure 2: The schematic overview of BiMamba Block.
-The BiMamba Block processes input sequences bidirectionally, capturing forward and reverse context through shared convolutional layers (Conv) and structured state machines (SSM).
-A gating mechanism integrates the outputs, followed by linear projection and nonlinearity layers, generating a context-aware representation for downstream tasks.
 3.1 Data Processing
@@
 The final normalized expression value is given by:
-$M_{ij}^{\text{norm}}=\frac{M_{ij}/\sum_{k=1}^{n}M_{ik}}{\text{t-digest}\{M_{kj}\mid M_{kj}>0\}}$
+$M_{ij}^{\text{norm}}=\frac{M_{ij}/\sum_{k=1}^{n}M_{ik}}{\text{t-digest}\{M_{kj}\mid M_{kj}>0\}}$ (1)
 Finally, we rank the genes within each cell in descending order based on their normalized expression values.
 This ranking approach highlights genes that distinguish cell states while deprioritizing universally high-expression housekeeping genes, ensuring they are assigned lower ranks in downstream analyses.
-$R_{i}=\operatorname{argsort}\left(-M_{ij}^{\text{norm}}\right),\quad\forall j$
+$R_{i}=\operatorname{argsort}\left(-M_{ij}^{\text{norm}}\right),\quad\forall j$ (2)
 3.2 Bi-Mamba Architecture
+Figure 2: The schematic overview of BiMamba Block.
+The BiMamba Block processes input sequences bidirectionally, capturing forward and reverse context through shared convolutional layers (Conv) and structured state machines (SSM).
+A gating mechanism integrates the outputs, followed by linear projection and nonlinearity layers, generating a context-aware representation for downstream tasks.
 State Space Models (SSMs) provide a powerful framework for modeling sequences by utilizing a latent state that evolves over time.
@@
 The dynamics of the system are described as:
-$h_{t}=Ah_{t-1}+Bx_{t},\quad y_{t}=Ch_{t}$
+$h_{t}=Ah_{t-1}+Bx_{t},\quad y_{t}=Ch_{t}$ (3)
 where $A,B,$ and $C$ are matrices defining how the input and state interact.
@@
 The convolutional view reformulates the output $y$ as:
-$y=x*K,\quad K=[CB,CAB,CA^{2}B,\dots]$
+$y=x*K,K=[CB,CAB,CA^{2}B,\dots]$ (4)
 where $K$ is a sequence-length-dependent kernel capturing temporal relationships.
@@
 In Mamba, the parameters $A,B,\Delta$, and $C$ are functions of the input, allowing the system to dynamically adjust to the sequence content. The updated equations are:
-$h_{t}=f_{A}(x_{t})h_{t-1}+f_{B}(x_{t})x_{t},\quad y_{t}=f_{C}(x_{t})h_{t}$
+$h_{t}=f_{A}(x_{t})h_{t-1}+f_{B}(x_{t})x_{t},\quad y_{t}=f_{C}(x_{t})h_{t}$ (5)
 where $f_{A},f_{B},$ and $f_{C}$ are learned transformations of the input $x_{t}$.
@@
 For each layer $l$, the outputs are:
-$h_{t}^{(l)}=f_{A}^{(l)}(s_{t})h_{t-1}^{(l)}+f_{B}^{(l)}(s_{t})s_{t}$
+$h_{t}^{(l)}=f_{A}^{(l)}(s_{t})h_{t-1}^{(l)}+f_{B}^{(l)}(s_{t})s_{t}$ (6)
 for the original sequence, and:
-$\tilde{h}_{t}^{(l)}=f_{A}^{(l)}(s_{t})\tilde{h}_{t-1}^{(l)}+f_{B}^{(l)}(s_{t})s_{t}$
+$\tilde{h}_{t}^{(l)}=f_{A}^{(l)}(s_{t})\tilde{h}_{t-1}^{(l)}+f_{B}^{(l)}(s_{t})s_{t}$ (7)
 for the reversed sequence $S_{\text{rev}}$, where $\tilde{h}_{t}^{(l)}$ represents the reversed latent state.
@@
 Instead of summing the outputs directly, Bi-Mamba introduces a learnable gating mechanism to balance the contributions of forward and backward information:
-$z_{t}^{(l)}=\sigma(W^{(l)}[h_{t}^{(l)},\tilde{h}_{t}^{(l)}])$
-$o_{t}^{(l)}=z_{t}^{(l)}\cdot h_{t}^{(l)}+(1-z_{t}^{(l)})\cdot\tilde{h}_{t}^{(l)}$
+$z_{t}^{(l)}=\sigma(W^{(l)}[h_{t}^{(l)},\tilde{h}_{t}^{(l)}])$ (8)
+$o_{t}^{(l)}=z_{t}^{(l)}\cdot h_{t}^{(l)}+(1-z_{t}^{(l)})\cdot\tilde{h}_{t}^{(l)}$ (9)
 where $\sigma$ is a sigmoid function, $W^{(l)}$ is a learnable weight matrix, and $o_{t}^{(l)}$ is the combined output for token $t$ at layer $l$.
 4. Stacking Layers: The Bi-Mamba process is repeated across multiple layers, with each layer refining the bidirectional representations:
-$S^{(l+1)}=\text{Bi-Mamba}(S^{(l)}),\quad S^{(0)}=S$
+$S^{(l+1)}=\text{Bi-Mamba}(S^{(l)}),\quad S^{(0)}=S$ (10)
 The final sequence representation $S^{(L)}$ encodes rich bidirectional dependencies, making it suitable for tasks requiring global context.
@@
 The loss function, referred to as $\mathcal{L}_{\text{lang}}$, is computed as the negative log-likelihood (NLL) of the true next gene token $g_{j}$ given the preceding tokens $\{g_{1},g_{2},\dots,g_{j-1}\}$, represented as:
-$\mathcal{L}_{\text{lang}}=-\frac{1}{M}\sum_{j=1}^{M}\log P(g_{j}\mid g_{1},g_{2},\dots,g_{j-1})$
+$\mathcal{L}_{\text{lang}}=-\frac{1}{M}\sum_{j=1}^{M}\log P(g_{j}\mid g_{1},g_{2},\dots,g_{j-1})$ (11)
 This loss function enables end-to-end training of the model by leveraging the sequential representation of gene expression data to predict each gene based on its preceding context within the same cell.
@@
 Gene Pathway Loss
-To capture the gene-gene relationships within biological pathways, we employ an InfoNCE loss function that enforces similarity among genes sharing a common pathway (Oord et al., 2018; Sharma & Xu, 2023; Gundogdu et al., 2022).
-Pathways represent functional groupings of genes that collectively contribute to specific biological processes (Zien et al., 2000; García-Campos et al., 2015).
+To capture the gene-gene relationships within biological pathways, we employ an InfoNCE loss function that enforces similarity among genes sharing a common pathway oord2018representation; sharma2023regenne; gundogdu2022integrating.
+Pathways represent functional groupings of genes that collectively contribute to specific biological processes zien2000analysis; garcia2015pathway.
 Genes within the same pathway are often functionally related and co-regulated, making their representations in the embedding space naturally similar.
@@
 Given a gene pair $(i,j)$, we compute the cosine similarity between their embeddings:
-$\text{sim}(i,j)=\frac{\mathbf{h}_{i}\cdot\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}$
+$\text{sim}(i,j)=\frac{\mathbf{h}_{i}\cdot\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}$ (12)
 where $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are the normalized embeddings of genes $i$ and $j$.
@@
 To emphasize the contrast between positive and negative pairs, we normalize the similarities using temperature scaling:
-$\tilde{\text{sim}}(i,j)=\frac{\text{sim}(i,j)}{\tau}$
+$\tilde{\text{sim}}(i,j)=\frac{\text{sim}(i,j)}{\tau}$ (13)
 where $\tau>0$ is a temperature hyperparameter. The InfoNCE loss for a batch of gene pairs is then formulated as:
-$\mathcal{L}_{\text{pathway}}=-\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\tilde{\text{sim}}(i,j))}{\sum_{k\in\mathcal{N}(i)}\exp(\tilde{\text{sim}}(i,k))}$
+$\mathcal{L}_{\text{pathway}}=-\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\tilde{\text{sim}}(i,j))}{\sum_{k\in\mathcal{N}(i)}\exp(\tilde{\text{sim}}(i,k))}$ (14)
 Here, $\mathcal{P}$ denotes the set of positive gene pairs (genes in the same pathway), and $\mathcal{N}(i)$ includes all gene pairs involving gene $i$ in the batch, including negatives.
@@
 The final pretraning loss is the weighted sum of the three loss:
-$\mathcal{L}=\mathcal{L}_{\text{lang}}+\gamma\mathcal{L}_{\text{pathway}}$
-We set the $\gamma$ value to 0.1 in our experiments based on validation results from sample datasets.
+$\mathcal{L}=\mathcal{L}_{\text{lang}}+\gamma\mathcal{L}_{\text{pathway}}$ (15)
 4 Experiments
@@
 Genes were then ranked based on their normalized expression levels within each cell, and the top 2,048 or 4,096 gene indices were selected as input features for the pretraining stage.
+The $\gamma$ value is set to 0.1 in our experiments based on validation results from sample datasets.
+The code and datasets are available at an anonymous address: https://anonymous.4open.science/r/GeneMamba-747D.
 Pretraining Configuration We configured GeneMamba with 24 Bi-Mamba layers, an inner dimension of 512, and a vocabulary size of 25,426, resulting in 65.74 million learnable parameters.
@@
 This highlights its robust capability in multi-batch integration and its biological relevance for single-cell data analysis.
-For further details, please refer to Figures 9 and 10.
+For further details, please refer to Figures 12 and 13.
 Table 1: Benchmark results of the models on multi-batch experiments with BATCH and Cell metrics
@@
 GeneMamba 0.6825 0.5342
+Datasets Models Acc Macro-F1
 Myeloid GeneFormer 0.6445 0.3600
@@
 4.3 Gene Rank Reconstruction
-Figure 4: Gene rank reconstruct results on PBMC12k dataset (a) Venn diagrams showing overlapping between input and output tokens in the pancreas dataset by three models: GeneMamba_U (unidirectional Mamba module as backbone), GeneFormer, GeneMamba (BiMamba module as backbone).
+Figure 4: Gene rank reconstruct results on PBMC12k dataset.
+(a) Venn diagrams showing overlapping between input and output tokens in the pancreas dataset by three models: GeneMamba_U (unidirectional Mamba module as backbone), GeneFormer, GeneMamba (BiMamba module as backbone).
 (b) Density plots showing input and output ranking in the pancreas dataset by three models: GeneMamba_U, GeneFormer, GeneMamba.
-The GeneMamba model demonstrates exceptional capability in reconstructing ranked gene orders, a critical requirement for single-cell analysis tasks where gene expression patterns reveal cellular states and transitions.
-The analysis begins by randomly selecting sample cells.
-From these cells, input gene tokens are extracted, and output tokens are generated using the GeneMamba model.
-Then the gene tokens are ranked based on their predicted likelihoods, with missing tokens filled by zeros to ensure sequence consistency.
-This approach preserves the structural integrity of the data and facilitates direct comparison between input and output rankings.
+The GeneMamba model exhibits strong capability in reconstructing ranked gene orders, a crucial requirement for single-cell analysis tasks where gene expression patterns reflect cellular states and transitions.
+This challenge has been highlighted in prior studies kedzierska2023assessing; boiarsky2023deep, and here we demonstrate the effectiveness of GeneMamba in addressing it.
+The analysis begins with the random selection of sample cells, from which input gene tokens are extracted.
+GeneMamba then generates output tokens, which are ranked according to their predicted likelihoods.
+To maintain sequence consistency, missing tokens are imputed with zeros.
+This strategy preserves the structural integrity of the data and enables a direct comparison between the input and output rankings.
+The consistency of GeneMamba’s predictions is validated through a Venn diagram Figure 4a, which illustrates a high degree of overlap between input and output tokens for the pancreas dataset.
+This overlap highlights the model’s ability to retain critical features of the input data, an essential characteristic for single-cell studies where the integrity of gene expression ranks directly influences downstream analyses.
+A density plot Figure 4b provides further evidence of rank fidelity.
+It demonstrates that lower-ranked input genes consistently yield lower-ranked outputs, indicating a linear relationship between input and output ranks.
+This observation underscores the model’s effectiveness in capturing patterns within the data and maintaining rank consistency.
 Table 3: Gene rank reconstruction performance on the PBMC12k dataset
@@
 GeneMamba 6 0.987 0.711
-The consistency of GeneMamba’s predictions is validated through a Venn diagram Figure 4a, which illustrates a high degree of overlap between input and output tokens for the pancreas dataset.
-This overlap highlights the model’s ability to retain critical features of the input data, an essential characteristic for single-cell studies where the integrity of gene expression ranks directly influences downstream analyses.
-A density plot Figure 4b provides further evidence of rank fidelity.
-It demonstrates that lower-ranked input genes consistently yield lower-ranked outputs, indicating a linear relationship between input and output ranks.
-This observation underscores the model’s effectiveness in capturing patterns within the data and maintaining rank consistency.
 Comparative analysis (Figure 4 from left to right) reveals the advantages of GeneMamba over other models: (1) The inclusion of the BiMamba module significantly enhances performance compared to the unidirectional module.
@@
 Additional results are provided in the Appendix Table 7.
-By reliably reconstructing ranked gene orders, GeneMamba offers a powerful solution for understanding complex gene expression patterns.
-Its ability to maintain rank integrity across varying data scales positions it as an invaluable tool for single-cell genomics.
-4.4 Gene Correlation Analysis
-Figure 5: Gene Correlation Analysis.
-(a) Distribution of consine similarity score for positive and negative gene pairs using Gene2Vec embeddings.
-(b) Distribution of pearson correlation score for positive and negative gene pairs using Gene2Vec embeddings.
-(c) Gene-gene topology comparison using adjacency matrices highlights shared and unique structures captured by GeneMamba, scGPT, and Gene2Vec, such as genes RALA, DECR1, and CBL1.
-GeneMamba exhibits exceptional capabilities in gene correlation analysis, distinguishing itself in capturing biologically meaningful patterns, maintaining raw data fidelity, and identifying unique model-specific topologies.
-A series of experiments evaluated GeneMamba’s performance in distinguishing gene relationships, preserving correlation structures, and classifying gene types.
-Differentiating Positive and Negative Gene Pairs In the first experiment, GeneMamba’s ability to differentiate between ”positive” and ”negative” gene pairs was assessed using Gene2Vec embeddings.
-Nearly 30,000 gene pairs, evenly distributed as positive (1) or negative (0), were analyzed for similarity scores using cosine similarity and Pearson correlation.
-GeneMamba demonstrated a significant separation between the mean similarity scores of positive and negative pairs (Figure 5 a, b).
-Compared to baseline models, GeneMamba achieved a clearer distinction, emphasizing its capacity to accurately represent gene relationships.
-Embedding Alignment To quantify discrimination power, we compared GeneMamba’s embeddings with those of scGPT and scFoundation.
-The ”distance” between positive and negative pair distributions was measured using Euclidean distance, KL divergence, and JS divergence.
-The result is shown in Appendix Figure 12.
-GeneMamba consistently outperformed the other models across all metrics, highlighting its superior ability to distinguish between gene pair types and reinforcing its robustness in modeling gene relationships.
-Gene-Gene Topology Analysis The ability to capture gene-gene topologies was analyzed using adjacency matrices derived from embeddings.
-For the PBMC12K dataset, distances between embeddings of randomly selected gene pairs (100 cells in our case) were calculated and binarized based on threshold value.
-Then we build the topology graph based on the adjacent matrix (Figure  5c) for each evaluated model.
-Notably, while identifying unique model-specific topologies, all three models captured biologically significant genes and their interactions, such as RALA, DECR1, and CBL1.
-This suggests that the foundation models share a similar representation space.
 5 Conclusion
-In this study, we present GeneMamba, a foundational model designed to advance single-cell analysis.
-By treating cells as sentences and genes as tokens, and with the BiMamba block and biological aware loss function, GeneMamba achieves a deeper understanding of cellular heterogeneity and gene relationships.
-This work underscores the potential of adapting large language models for biological data analysis, bridging computational innovation with biological insights.
-Future directions include enhancing the model with mechanisms like cross-attention for improved transfer learning, scaling to even larger datasets, and creating accessible tools for the research community.
-Code Availability
-Datasets and code are available at the github address: https://github.com/MineSelf2016/GeneMamba
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. Our work has potential societal implications, but none warrant specific emphasis here.
+We present GeneMamba, a scalable model for single-cell analysis that treats cells as sentences and genes as tokens.
+Using the BiMamba block and biologically informed losses, GeneMamba captures gene relationships and cell diversity effectively.
+It performs well across key tasks and offers a strong foundation for future single-cell studies.
+Broader Impacts and Limitations.
+GeneMamba can help researchers analyze large-scale single-cell data, offering insights into gene function and cell behavior.
+This has potential benefits in areas like disease research and drug development.
+However, the model may struggle with rare cell types or subtle signals from low-expression genes.
+It also requires significant computing resources for pretraining, which could limit broader use.
+Future work could focus on improving efficiency and expanding to more flexible architectures.
 Appendix
-Appendix A Dataset Construction
+Appendix A Dataset Construction and Setup
 We constructed our pretraining dataset manually, which is sourced from the cellxgene database, acquiring single-cell RNA sequencing (scRNA-seq) data in raw count matrix format alongside the corresponding metadata.
@@
 The original Myeloid dataset is highly imbalanced.
-To better evaluate the models’ performance, we excluded rare cell types (proportion $<$ 0.05) to create the Myeloid_b dataset.
+To better evaluate the models’ performance, we excluded rare cell types (proportion << 0.05) to create the Myeloid_b dataset.
 Figure 5 illustrates the cell type distribution in both the original and modified Myeloid datasets.
@@
 B.2 Downstream Tasks
-We standardize the downstream datasets by converting the batch key to ”batch” and the label key to ”celltype.” We then use a manual stratified split to divide the dataset into train/test with a 0.9/0.1 ratio to avoid missing rare classes in the test set.
+We standardize the downstream datasets by converting the batch key to "batch" and the label key to "celltype." We then use a manual stratified split to divide the dataset into train/test with a 0.9/0.1 ratio to avoid missing rare classes in the test set.
 For example, in the COVID dataset, using ‘train_test_split()‘ directly would cause the 22nd class to be missing, as it contains only 3 samples, and a 0.3 test sample rate would result in zero samples, causing errors during AUC-ROC calculation.
-The split column is labeled ”partition,” where ”train” is for training and ”test” is for testing.
+The split column is labeled "partition," where "train" is for training and "test" is for testing.
 Multi-batch integration Task
@@
 Gene Rank Reconstruction
-Figure 7: Gene Rank Reconstruct framework of GeneMamba Gene Imputation framework evaluating rank reconstruction by predicting (n+1)-th tokens iteratively.
+Figure 6: Gene Rank Reconstruct framework of GeneMamba
 In the figure 6, the framework of doing the gene rank reconstruction experiment.
@@
 Gene Distribution Alignment
-In this experiment, we evaluate the GeneMamba’s performance to distinguish between positive pairs and negative pairs, the reference data is collected in the Gene2Vec paper (Du et al., 2019).
+In this experiment, we evaluate the GeneMamba’s performance to distinguish between positive pairs and negative pairs, the reference data is collected in the Gene2Vec paper [du2019gene2vec].
 And we generate the embeddings using different models for positive gene pairs and negative gene pairs, then we measure the embeddings similarity of each pair.
@@
 To evaluate the performance of our proposed approach, we compare it with several state-of-the-art baseline methods. Below, we briefly describe each method:
-  • GeneFormer (Theodoris et al., 2023): A transformer-based model specifically designed for single-cell analysis.
+  • GeneFormer [theodoris2023transfer]: A transformer-based model specifically designed for single-cell analysis.
 It uses a gene-token representation to capture complex gene-gene relationships and downstream cellular heterogeneity.
-  • scGPT (Cui et al., 2024a): A generative pre-trained transformer adapted for single-cell data analysis.
+  • scGPT [cui2024scgpt]: A generative pre-trained transformer adapted for single-cell data analysis.
 scGPT utilizes a pretraining-finetuning paradigm to handle various single-cell tasks, including cell-type classification and clustering.
-  • scFoundation (Hao et al., 2024): A foundational model for single-cell data analysis that incorporates self-supervised learning techniques to generate robust embeddings for single-cell profiles.
+  • scFoundation [hao2024large]: A foundational model for single-cell data analysis that incorporates self-supervised learning techniques to generate robust embeddings for single-cell profiles.
 It excels in tasks requiring integration across batches and datasets.
-  • scBERT (Yang et al., 2022): A BERT-like architecture adapted for single-cell data.
+  • scBERT [yang2022scbert]: A BERT-like architecture adapted for single-cell data.
 scBERT leverages bidirectional contextual embeddings to model complex interactions between genes within cells, making it suitable for both classification and clustering tasks.
-  • Harmony (Korsunsky et al., 2019): A batch-effect correction method that aligns single-cell datasets from different conditions or experiments.
+  • Harmony [korsunsky2019fast]: A batch-effect correction method that aligns single-cell datasets from different conditions or experiments.
 This method is often used as a preprocessing step to ensure integration and compatibility across datasets before applying downstream machine learning methods.
@@
   • scFoundation: Since the scFoundation model does not provide a direct method for obtaining gene embeddings, we apply mean pooling to the generated context-aware embeddings of scFoundation on downstream datasets.
-Appendix C More Related Work
-C.1 Discretization of Gene Expression
+Appendix C Gene Correlation Analysis
+Figure 7: Gene Correlation Analysis.
+(a) Distribution of consine similarity score for positive and negative gene pairs using Gene2Vec embeddings.
+(b) Distribution of pearson correlation score for positive and negative gene pairs using Gene2Vec embeddings.
+(c) Gene-gene topology comparison using adjacency matrices highlights shared and unique structures captured by GeneMamba, scGPT, and Gene2Vec, such as genes RALA, DECR1, and CBL1.
+GeneMamba exhibits exceptional capabilities in gene correlation analysis, distinguishing itself in capturing biologically meaningful patterns, maintaining raw data fidelity, and identifying unique model-specific topologies.
+A series of experiments evaluated GeneMamba’s performance in distinguishing gene relationships, preserving correlation structures, and classifying gene types.
+Differentiating Positive and Negative Gene Pairs In the first experiment, GeneMamba’s ability to differentiate between "positive" and "negative" gene pairs was assessed using Gene2Vec embeddings.
+Nearly 30,000 gene pairs, evenly distributed as positive (1) or negative (0), were analyzed for similarity scores using cosine similarity and Pearson correlation.
+GeneMamba demonstrated a significant separation between the mean similarity scores of positive and negative pairs (Figure 7 a, b).
+Compared to baseline models, GeneMamba achieved a clearer distinction, emphasizing its capacity to accurately represent gene relationships.
+Embedding Alignment To quantify discrimination power, we compared GeneMamba’s embeddings with those of scGPT and scFoundation.
+The "distance" between positive and negative pair distributions was measured using Euclidean distance, KL divergence, and JS divergence.
+The result is shown in Appendix Figure 15.
+GeneMamba consistently outperformed the other models across all metrics, highlighting its superior ability to distinguish between gene pair types and reinforcing its robustness in modeling gene relationships.
+Gene-Gene Topology Analysis The ability to capture gene-gene topologies was analyzed using adjacency matrices derived from embeddings.
+For the PBMC12K dataset, distances between embeddings of randomly selected gene pairs (100 cells in our case) were calculated and binarized based on threshold value.
+Then we build the topology graph based on the adjacent matrix (Figure  7c) for each evaluated model.
+Notably, while identifying unique model-specific topologies, all three models captured biologically significant genes and their interactions, such as RALA, DECR1, and CBL1.
+This suggests that the foundation models share a similar representation space.
+Appendix D Perturbation Analysis
+To more comprehensively evaluate the generalizability and biological relevance of GeneMamba, we conducted a perturbation prediction study using protocols established by scGPT [cui2024scgpt].
+Perturbation studies are critical for assessing a model’s ability to simulate cellular responses to genetic interventions—a fundamental requirement in functional genomics and drug discovery.
+We employed the Perturb-seq dataset from Norman [norman2019exploring], which include two-gene combinatorial perturbations.
+Specifically, we compared Mamba-based GEARS embeddings with the original GEARS model (baseline), following the benchmarking criteria previously proposed.
+Figure 8: Mean Squared Error of DE analysis
+Since single-cell-level ground truth was unavailable, we evaluated performance using the averaged Mean Squared Error (MSE) of the top 20 differentially expressed (DE) genes before and after perturbation, and the result is shown in Figure 8.
+GeneMamba achieved lower MSE across these genes compared to the baseline, suggesting improved fidelity in modeling transcriptional changes.
+Additionally, we quantified prediction accuracy by measuring the proportion of top 20 DE genes whose predicted expression fell within the 45–55% quantile range of the true expression distribution, where the Mamba-based model again showed superior alignment with ground truth.
+Figure 9: ETS2 + CEBPE perturbation result
+Figure 10: Magnitude scores computed for all test perturbing combinations on the Norman dataset.
+Furthermore, in a representative combinatorial perturbation case (ETS2 + CEBPE), as shown in Figure 9, GeneMamba-generated predictions (red) closely matched the ground truth (purple) and outperformed the baseline (blue) in capturing post-perturbation expression shifts.
+Correlation analysis between predicted and ground-truth magnitude scores across all test perturbations yielded a higher Pearson coefficient (r = 0.5686, p < 1e-6) for GeneMamba, compared to the baseline GEARS (r = 0.4117), as shown in Figure 10.
+Lastly, the identification of suppressor and synergistic gene interactions revealed broader and more accurate overlap with the verified perturbation set using GeneMamba.
+These findings underscore the model’s enhanced interpretability and robustness under gene perturbation scenarios.
+Appendix E More Related Work
+E.1 Discretization of Gene Expression
 The transformation of gene expression levels into machine-readable tokens remains a critical aspect of single-cell modeling, with emerging methods aiming to balance biological relevance and computational efficiency.
@@
 Genes falling within the same interval are represented identically.
-This method, however, often leads to significant information loss, as most genes with low expression levels cluster within the same bin (e.g., the [0,1) bin (Boiarsky et al., 2023)), masking subtle but biologically relevant differences.
+This method, however, often leads to significant information loss, as most genes with low expression levels cluster within the same bin (e.g., the [0,1) bin [boiarsky2023deep]), masking subtle but biologically relevant differences.
 Adaptive Binning: Adaptive binning, employed by scGPT, dynamically adjusts bin sizes such that each interval represents an equal proportion of expressed genes within a single cell.
 While this approach aims to better capture the distribution of gene expression, it still faces challenges related to information loss, particularly for genes with extreme expression values or biological significance.
-Value Projection Strategy: The value projection strategy, first introduced by TOSICA (Chen et al., 2023), has since been adopted by scCLIP (Xiong et al., 2023), SpaFormer (Wen et al., 2023b), CellLM (Zhao et al., 2023), and CellPLM (Wen et al., 2023a).
+Value Projection Strategy: The value projection strategy, first introduced by TOSICA [chen2023transformer], has since been adopted by scCLIP [xiong2023scclip], SpaFormer [wen2023single], CellLM [zhao2023large], and CellPLM [wen2023cellplm].
 In this approach, continuous gene expression values are directly projected into high-dimensional embeddings through linear or non-linear transformations.
@@
 Additionally, continuous embeddings may be more sensitive to noise and batch effects, necessitating robust preprocessing steps.
-Rank-Based Strategy: The rank-based strategy was first introduced in iSEEEK (Shen et al., 2022) for single-cell transformers.
+Rank-Based Strategy: The rank-based strategy was first introduced in iSEEEK [shen2022universal] for single-cell transformers.
 This approach ranks gene expression values within each cell and retains only the top-ranked genes.
@@
 It also aligns well with the biological prioritization of highly expressed genes in specific cell states or conditions.
-C.2 Model Architectures for Single-Cell Analysis
-Recent advancements in model architectures have addressed the computational challenges of transformer-based models in single-cell RNA sequencing (scRNA-seq) analysis, with notable innovations including scTCA (Yu et al., 2024), a hybrid Transformer-CNN architecture designed for imputation and denoising of single-cell read counts to enhance downstream analyses; scHyena (Oh et al., 2023), a foundation model for full-length scRNA-seq analysis in brain tissues that uses a novel transformer architecture to process data without information loss, improving tasks like cell type classification and data imputation; xTrimoGene (Gong et al., 2024b), an efficient and scalable representation learner employing an asymmetric encoder-decoder transformer architecture to reduce computational requirements while maintaining high accuracy in tasks such as cell type annotation and drug combination prediction; scTransSort (Jiao et al., 2023), a transformer-based model for intelligent cell type annotation that leverages self-attention mechanisms to extract features from scRNA-seq data, achieving high accuracy and performance; scGraphformer (Fan et al., 2024), which employs transformer-based graph neural networks to provide accurate and scalable cell type annotations, uncovering cellular heterogeneity and interactions in scRNA-seq data; scGFT (Nouri, 2025), a train-free, cell-centric generative model adept at synthesizing scRNA-seq data to enhance data augmentation and analysis; White-Box Diffusion Transformer(Cui et al., 2024b), a hybrid model combining diffusion models and white-box transformers to generate synthetic and biologically plausible scRNA-seq data, improving training efficiency and resource utilization; STGRNS (Xu et al., 2023), an interpretable transformer-based method for inferring gene regulatory networks from scRNA-seq data to enhance understanding of gene interactions; scRDiT (Dong et al., 2024), a generative approach using diffusion transformers to create virtual scRNA-seq data, facilitating the study of gene expression dynamics; and TransformerST (Zhao et al., 2024a), an unsupervised model based on the transformer architecture for super-resolution in spatial transcriptomics, improving spatial gene expression analysis.
+E.2 Model Architectures for Single-Cell Analysis
+Recent advancements in model architectures have addressed the computational challenges of transformer-based models in single-cell RNA sequencing (scRNA-seq) analysis, with notable innovations including scTCA [scTCA], a hybrid Transformer-CNN architecture designed for imputation and denoising of single-cell read counts to enhance downstream analyses; scHyena [scHyena], a foundation model for full-length scRNA-seq analysis in brain tissues that uses a novel transformer architecture to process data without information loss, improving tasks like cell type classification and data imputation; xTrimoGene [gong2024xtrimogene], an efficient and scalable representation learner employing an asymmetric encoder-decoder transformer architecture to reduce computational requirements while maintaining high accuracy in tasks such as cell type annotation and drug combination prediction; scTransSort [scTransSort], a transformer-based model for intelligent cell type annotation that leverages self-attention mechanisms to extract features from scRNA-seq data, achieving high accuracy and performance; scGraphformer [scGraphformer], which employs transformer-based graph neural networks to provide accurate and scalable cell type annotations, uncovering cellular heterogeneity and interactions in scRNA-seq data; scGFT [scGFT], a train-free, cell-centric generative model adept at synthesizing scRNA-seq data to enhance data augmentation and analysis; White-Box Diffusion Transformer[WhiteBoxDiffusionTransformer], a hybrid model combining diffusion models and white-box transformers to generate synthetic and biologically plausible scRNA-seq data, improving training efficiency and resource utilization; STGRNS [STGRNS], an interpretable transformer-based method for inferring gene regulatory networks from scRNA-seq data to enhance understanding of gene interactions; scRDiT [scRDiT], a generative approach using diffusion transformers to create virtual scRNA-seq data, facilitating the study of gene expression dynamics; and TransformerST [TransformerST], an unsupervised model based on the transformer architecture for super-resolution in spatial transcriptomics, improving spatial gene expression analysis.
 These models collectively advance the field by addressing key challenges in scRNA-seq data processing and interpretation.
-Traditional single-cell large language models (LLMs) are predominantly based on transformer architectures (Boiarsky et al., 2023; White & Harris, 2024).
+Traditional single-cell large language models (LLMs) are predominantly based on transformer architectures [boiarsky2023deep, white2024comprehensive].
 While transformers have demonstrated exceptional performance in various single-cell transcriptomics tasks, their high computational and memory requirements make training large models time-intensive and resource-heavy.
@@
 To alleviate these challenges, some transformer-based models have incorporated optimizations to improve efficiency.
-For example, scBERT integrates Performer (Choromanski et al., 2020) to approximate attention calculations, while scGPT employs Flash Attention (Dao et al., 2022) to accelerate training by reducing memory overhead.
+For example, scBERT integrates Performer [choromanski2020rethinking] to approximate attention calculations, while scGPT employs Flash Attention [dao2022flashattention] to accelerate training by reducing memory overhead.
 Despite these advancements, the overall time required to train such models remains substantial, especially for datasets with millions of cells and thousands of genes.
@@
 This represents a significant departure from traditional transformer-based methods, offering a scalable and efficient solution for the high-dimensional, high-throughput nature of single-cell transcriptomics data.
-Appendix D Discussion of Scalability
+Appendix F Discussion of Scalability
 Scalability is an essential feature for models designed to handle large-scale single-cell omics datasets.
@@
 Additionally, GeneMamba demonstrates its capability to handle longer sequence lengths (e.g., 4096 tokens) with a linear increase in training time (from 5.5 hours to 11 hours), showcasing its scalability for more input gene tokens of single-cell data.
-Mamba Layers Hidden Size Seq Length Number of Params Training Time (hours/m)
+Table 5: Training configurations for the GeneMamba models, demonstrating scalability across layers, hidden sizes, and sequence lengths.
+The data highlights the increasing number of parameters and training time as model complexity grows, with GeneMamba showing efficient handling of longer sequence lengths (e.g., 4096) compared to traditional Mamba setups.
+Mamba Layers Hidden Size Seq Length Param Size Training Time (h/million)
 24 512 2048 65.74M 5.5
@@
 24 512 4096 65.74M 11
-Table 5: Training configurations for the GeneMamba models, demonstrating scalability across layers, hidden sizes, and sequence lengths.
-The data highlights the increasing number of parameters and training time as model complexity grows, with GeneMamba showing efficient handling of longer sequence lengths (e.g., 4096) compared to traditional Mamba setups.
 Computational Efficiency Comparison
@@
 This efficiency becomes even more pronounced at longer sequence lengths.
-Model Backbone Parameters Forward + Backward FLOPs/sample Seq Length Training Time (seconds/sample)
+These results confirm that our conclusions remain valid: FlashAttention only optimizes memory access but keeps $O(L^{2})$ scaling, whereas GeneMamba achieves true linear-time scaling and much lower computational cost for long sequences.
+Table 6: Comparison of computational efficiency across Transformer, FlashAttention, and GeneMamba.
+Model Backbone Parameters FLOPs/sample Seq Length Training Time (s/sample)
 Transformer 130M 2.85E+12 2048 0.2227
@@
 Transformer 130M 2.02E+13 8192 1.8904
+FlashAttention 130M 2.62E+12 2048 0.2087
+FlashAttention 130M 7.03E+12 4096 0.5899
+FlashAttention 130M 1.96E+13 8192 1.8342
 GeneMamba 130M 1.58E+12 2048 0.1284
@@
 GeneMamba 130M 6.31E+12 8192 0.4885
-Table 6: Computational efficiency comparison of Transformer and GeneMamba backbones at different input sequence lengths.
-GeneMamba significantly reduces computational cost while maintaining high efficiency.
 Theoretical Computation and Demonstration
 Using the 2048-sequence-length GeneMamba model as an example, the total FLOPs required for processing 2 million cells can be calculated as:
-Total FLOPs $\text{Total FLOPs required}=1.58\times 10^{12}\times 2\times 10^{6}=3.16\times 10^{18}\text{ FLOPs.}$.
+$\text{Total FLOPs required}=1.58\times 10^{12}\times 2\times 10^{6}=3.16\times 10^{18}\text{ FLOPs.}$
 Given that the A100 80GB GPU has a peak performance of 19.5 TFLOPS for single-precision (FP32) and 312 TFLOPS for half-precision (FP16), the theoretical training time is computed as follows:
 For FP32:
-$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{19.5\times 10^{12}\text{ FLOPs/sec}}\approx 1.621\times 10^{5}\text{ seconds}=45\text{ hours.}$.
+$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{19.5\times 10^{12}\text{ FLOPs/sec}}\approx 1.621\times 10^{5}\text{ seconds}=45\text{ hours.}$
 For FP16:
-$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{312\times 10^{12}\text{ FLOPs/sec}}\approx 1.013\times 10^{4}\text{ seconds}=2.8\text{ hours.}$.
+$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{312\times 10^{12}\text{ FLOPs/sec}}\approx 1.013\times 10^{4}\text{ seconds}=2.8\text{ hours.}$
 These results demonstrate the substantial efficiency improvements provided by GeneMamba, especially when leveraging FP16 precision, which reduces training time from 45 hours to just 2.8 hours.
@@
 As shown in Tables 5 and 6, its scalability and efficiency significantly outperform traditional Transformer architectures, providing a robust framework for computational biology and related applications.
-Appendix E Extended Figures
+Appendix G Extended Figures
 Figure 11: Results of cell type annotation. Bar plot of the classification metrics across various datasets of GeneMamba.
@@
 The embeddings generated by GeneMamba are more similar to Gene2Vec embeddings compared to randomly generated embeddings.
-Appendix F Extended Tables
+Appendix H Extended Tables
 Table 7: Gene Rank Reconstruction Performance comparison of GeneMamba, GeneFormer, and GeneMamba_S models across datasets.
@@
 GeneMamba 12 0.998 0.844
+We present the complete baseline results directly below, including both Accuracy and Macro-F1 scores, along with the detailed experimental configurations used to obtain them.
+Table 8: Baseline results comparing HVG, Seurat, scVI, scANVI, and GeneMamba. Bold numbers denote best performance.
+Dataset Metric HVG Seurat scVI scANVI GeneMamba
+hPancreas Accuracy 0.9770 0.9668 0.9597 0.9662 0.9713
+Macro-F1 0.7029 0.7346 0.6631 0.7023 0.7710
+MS Accuracy 0.7055 0.8600 0.7790 0.8675 0.6825
+Macro-F1 0.5403 0.7335 0.6357 0.7424 0.5342
+Myeloid Accuracy 0.6052 0.6571 0.6029 0.6332 0.6607
+Macro-F1 0.3334 0.3627 0.3330 0.3400 0.3650
+Myeloid_b Accuracy 0.8942 0.9018 0.8929 0.9231 0.9603
+Macro-F1 0.8825 0.8930 0.8780 0.9147 0.9235
+Experimental Setup
+  • HVG (Highly Variable Genes): We selected the top 2,000 highly variable genes per dataset to reduce dimensionality while retaining biologically informative variation.
+  • Seurat (R v4.4.0):
+    1. 1.
+      Standard PCA for dimensionality reduction.
+    2. 2.
+      Random Forest as backbone module.
+    3. 3.
+      Clustering resolution was tuned to align with reference labels for fair comparison.
+  • scVI / scANVI (scvi-tools==1.1.0):
+    1. 1.
+      Latent space dimension: 32
+    2. 2.
+      Classification head: 2-layer MLP
+    3. 3.
+      Activation: ReLU
+    4. 4.
+      scANVI initialized from scVI embeddings with default semi-supervised ratio.
+  • Training:
+    1. 1.
+      100 epochs
+    2. 2.
+      Batch size: 512
+    3. 3.
+      Adam optimizer, learning rate: 1e-3
+    4. 4.
+      Early stopping was monitored on validation loss.
+Analysis and Observations
+  1. 1.
+    Overall performance: GeneMamba achieves the best performance on 3 of the 4 datasets and shows particularly strong Macro-F1 scores, indicating superior recognition of minority or rare cell populations.
+  2. 2.
+    Macro-F1 vs Accuracy:
+    • scANVI achieves competitive or slightly higher Accuracy on some datasets (e.g., MS) but falls short in Macro-F1, implying its predictions are skewed toward majority classes.
+    • GeneMamba’s higher Macro-F1 demonstrates balanced prediction across diverse cell types.
+  3. 3.
+    Biological implications: Balanced performance is crucial in single-cell analysis, as rare populations often carry key biological insights.
+GeneMamba’s results reflect its strength in capturing both majority and minority cell behaviors.
+Below, we present the results across four benchmark datasets, comparing GeneMamba with six strong baselines:
+Cell Type Annotation Performance (Expanded)
+Table 9: Expanded benchmarks comparing GeneMamba with six strong baselines. Bold values denote the best performance.
+Dataset Metric GeneFormer scGPT scFoundation scBERT scMulan GeneMamba
+hPancreas Accuracy 0.9665 0.9710 0.9602 0.9594 0.9501 0.9713
+Macro-F1 0.7450 0.7632 0.7101 0.7381 0.7012 0.7710
+MS Accuracy 0.7650 0.8471 0.7763 0.8054 0.7888 0.6825
+Macro-F1 0.6220 0.6630 0.6812 0.6412 0.6450 0.5342
+Myeloid Accuracy 0.6445 0.6341 0.6446 0.6527 0.6483 0.6607
+Macro-F1 0.3600 0.3562 0.3646 0.3482 0.3590 0.3650
+Myeloid_b Accuracy 0.9540 0.9421 0.9574 0.9451 0.9487 0.9603
+Macro-F1 0.9380 0.9434 0.9569 0.9261 0.9340 0.9235
+Interpretation: GeneMamba achieves the best performance in both Accuracy and Macro-F1 on hPancreas and Myeloid, two biologically diverse datasets.
+Even in more challenging or imbalanced datasets such as MS and Myeloid_b, it remains competitive.
+These additional results provide stronger empirical evidence that GeneMamba is a robust and effective foundation model, standing out among both classical transformer-based and modern state space models.
+Note: Results for scBERT and scMulan were obtained by fine-tuning their publicly released pretrained models on our standardized downstream annotation datasets, using the same MLP head, learning rate, and stratified train/test split as described in Appendix B.2.

### v3

@@
 1 Introduction
-Single-cell RNA sequencing (scRNA-seq) has fundamentally transformed our ability to study cellular heterogeneity and dynamics by enabling high-resolution profiling of gene expression at the individual cell level korsunsky2019fast; theodoris2023transfer.
-This technology allows researchers to dissect complex cellular compositions within tissues, trace differentiation trajectories, and identify rare cell populations that remain undetectable with bulk RNA sequencing shen2023generative; wen2023single; deconinck2021recent.
-The rich transcriptomic insights provided by scRNA-seq have catalyzed significant advancements in developmental biology, disease modeling, and drug discovery gu2023mamba; yang2024genecompass; li2024single.
-Despite these advantages, the computational analysis of scRNA-seq data presents formidable challenges due to its high dimensionality, inherent sparsity, and technical noise du2019gene2vec; zhao2023large.
+Single-cell RNA sequencing (scRNA-seq) has fundamentally transformed our ability to study cellular heterogeneity and dynamics by enabling high-resolution profiling of gene expression at the individual cell level Korsunsky et al.
+(2019); Theodoris et al.
+(2023).
+This technology allows researchers to dissect complex cellular compositions within tissues, trace differentiation trajectories, and identify rare cell populations that remain undetectable with bulk RNA sequencing Shen et al.
+(2023); Wen et al.
+(2023b); Deconinck et al.
+(2021).
+The rich transcriptomic insights provided by scRNA-seq have catalyzed significant advancements in developmental biology, disease modeling, and drug discovery Gu and Dao (2023); Yang et al.
+(2024); Li and Wang (2024).
+Despite these advantages, the computational analysis of scRNA-seq data presents formidable challenges due to its high dimensionality, inherent sparsity, and technical noise Du et al.
+(2019); Zhao et al.
+(2023).
 Addressing these challenges requires robust computational models capable of capturing biologically meaningful patterns while efficiently handling the scale and variability of single-cell data.
-Transformer-based architectures, such as scBERT yang2022scbert and scGPT cui2024scgpt, have emerged as powerful tools in this domain, demonstrating strong performance in tasks such as cell type classification, gene expression imputation, and differential expression analysis bian2024scmulan; wen2023cellplm; szalata2024transformers.
+Transformer-based architectures, such as scBERT Yang et al.
+(2022) and scGPT Cui et al.
+(2024a), have emerged as powerful tools in this domain, demonstrating strong performance in tasks such as cell type classification, gene expression imputation, and differential expression analysis Bian et al.
+(2024); Wen et al.
+(2023a); Szałata et al.
+(2024).
 However, transformers exhibit critical limitations when applied to scRNA-seq data.
-The quadratic complexity of their self-attention mechanism constrains their scalability for long sequences, which are characteristic of single-cell transcriptomes vaswani2017attention; choromanski2020rethinking.
-Moreover, transformers often struggle to effectively capture long-range dependencies, which are essential for modeling gene regulatory interactions and cell state transitions dao2022flashattention; chen2023transformer.
-These limitations have driven the exploration of alternative architectures, among which state space models (SSMs) have emerged as promising solutions, offering improved efficiency and scalability for processing long sequences gu2023mamba; mamba2.
+The quadratic complexity of their self-attention mechanism constrains their scalability for long sequences, which are characteristic of single-cell transcriptomes Vaswani et al.
+(2017); Choromanski et al.
+(2020).
+Moreover, transformers often struggle to effectively capture long-range dependencies, which are essential for modeling gene regulatory interactions and cell state transitions Dao et al.
+(2022); Chen et al.
+(2023).
+These limitations have driven the exploration of alternative architectures, among which state space models (SSMs) have emerged as promising solutions, offering improved efficiency and scalability for processing long sequences Gu and Dao (2023); Dao and Gu (2024).
 In this study, we introduce GeneMamba, a novel state space model designed to efficiently train large-scale cell models on scRNA-seq data.
-SSM-based models have demonstrated competitive or superior performance compared to transformers while significantly reducing computational overhead shen2022universal; gong2024xtrimogene; ryu2023integration.
-Building upon this foundation, GeneMamba incorporates bidirectional computation, enhancing its ability to capture both upstream and downstream contextual dependencies, thereby improving its performance in tasks requiring global contextual awareness liu2024bidirectional; liang2024bi.
+SSM-based models have demonstrated competitive or superior performance compared to transformers while significantly reducing computational overhead Shen et al.
+(2022); Gong et al.
+(2024); Ryu et al.
+(2023).
+Building upon this foundation, GeneMamba incorporates bidirectional computation, enhancing its ability to capture both upstream and downstream contextual dependencies, thereby improving its performance in tasks requiring global contextual awareness Liu et al.
+(2024); Liang et al.
+(2024).
 We validate GeneMamba across a diverse set of applications, including multi-batch integration, cell type annotation, and gene-gene correlation analysis.
-Our experimental results demonstrate substantial improvements in computational efficiency and predictive performance compared to existing models hao2024large; sun2024bidirectional; petegrosso2020machine.
+Our experimental results demonstrate substantial improvements in computational efficiency and predictive performance compared to existing models Hao et al.
+(2024); Sun et al.
+(2024); Petegrosso et al.
+(2020).
 Furthermore, to assess its scalability, we conduct a comprehensive performance analysis on various variations of GeneMamba (Appendix F), highlighting its advantages in real-world biological applications.
@@
 Existing methods employ various tokenization strategies, each with distinct advantages and limitations.
-Bin-based discretization, as used by scBERT yang2022scbert, scGPT cui2024scgpt, and scMulan bian2024scmulan, groups expression values into predefined bins.
+Bin-based discretization, as used by scBERT Yang et al.
+(2022), scGPT Cui et al.
+(2024a), and scMulan Bian et al.
+(2024), groups expression values into predefined bins.
 This approach preserves absolute value distributions and simplifies sequence modeling, but may introduce information loss, particularly for genes with subtle but biologically significant expression differences.
 Additionally, binning can be sensitive to parameter selection, affecting downstream results.
-Value projectionszalata2024transformers, adopted by scFoundation hao2024large and its backbone model xTrimoGene gong2024xtrimogene, projects gene expression values into continuous embeddings rather than discrete categories.
+Value projectionSzałata et al.
+(2024), adopted by scFoundation Hao et al.
+(2024) and its backbone model xTrimoGene Gong et al.
+(2024), projects gene expression values into continuous embeddings rather than discrete categories.
 This method maintains full data resolution by applying a linear transformation to the gene expression vector, which is then combined with gene-specific embeddings.
 However, the use of continuous embeddings diverges from traditional tokenization strategies in NLP-based transformers, and its impact on model performance remains an open question.
-Rank-based discretization, utilized by Geneformer theodoris2023transfer, GeneCompass yang2024genecompass, tGPTshen2023generative and LangCell zhao2024langcell, transforms gene expression values into ordinal rankings.
+Rank-based discretization, utilized by Geneformer Theodoris et al.
+(2023), GeneCompass Yang et al.
+(2024), tGPTShen et al.
+(2023) and LangCell Zhao et al.
+(2024b), transforms gene expression values into ordinal rankings.
 This approach effectively captures relative expression levels and is more robust to batch effects and noise.
@@
 These models have shown promising results in various tasks, including cell-type annotation, batch integration, and multiomics analysis.
-However, transformervaswani2017attention has inherent limitations.
+However, transformerVaswani et al.
+(2017) has inherent limitations.
 The most significant one is their quadratic computational complexity with respect to sequence length, which makes them less feasible for long sequences typical of scRNA-seq data.
@@
 Additionally, transformer often struggles with capturing long-range dependencies in sequences, which is crucial for understanding gene regulatory networks and cell state transitions.
-Mambagu2023mamba has been introduced to address the transformer efficiency problem.
-Bidirectional Mamba(Bi-Mamba) has recently been used to solve long-sequence problems in tasks such as feature extractionsun2024bidirectional, sequential recommendationliu2024bidirectional, and time series forecastingliang2024bi.
+MambaGu and Dao (2023) has been introduced to address the transformer efficiency problem.
+Bidirectional Mamba(Bi-Mamba) has recently been used to solve long-sequence problems in tasks such as feature extractionSun et al.
+(2024), sequential recommendationLiu et al.
+(2024), and time series forecastingLiang et al.
+(2024).
 Bi-Mamba is designed to efficiently process ultra-long sequences with linear computational complexity, offering a significant reduction in memory and computation requirements compared to transformer.
@@
 Gene Pathway Loss
-To capture the gene-gene relationships within biological pathways, we employ an InfoNCE loss function that enforces similarity among genes sharing a common pathway oord2018representation; sharma2023regenne; gundogdu2022integrating.
-Pathways represent functional groupings of genes that collectively contribute to specific biological processes zien2000analysis; garcia2015pathway.
+To capture the gene-gene relationships within biological pathways, we employ an InfoNCE loss function that enforces similarity among genes sharing a common pathway Oord et al.
+(2018); Sharma and Xu (2023); Gundogdu et al.
+(2022).
+Pathways represent functional groupings of genes that collectively contribute to specific biological processes Zien et al.
+(2000); García-Campos et al.
+(2015).
 Genes within the same pathway are often functionally related and co-regulated, making their representations in the embedding space naturally similar.
@@
 The GeneMamba model exhibits strong capability in reconstructing ranked gene orders, a crucial requirement for single-cell analysis tasks where gene expression patterns reflect cellular states and transitions.
-This challenge has been highlighted in prior studies kedzierska2023assessing; boiarsky2023deep, and here we demonstrate the effectiveness of GeneMamba in addressing it.
+This challenge has been highlighted in prior studies Kedzierska et al.
+(2023); Boiarsky et al.
+(2023), and here we demonstrate the effectiveness of GeneMamba in addressing it.
 The analysis begins with the random selection of sample cells, from which input gene tokens are extracted.
@@
 Gene Distribution Alignment
-In this experiment, we evaluate the GeneMamba’s performance to distinguish between positive pairs and negative pairs, the reference data is collected in the Gene2Vec paper [du2019gene2vec].
+In this experiment, we evaluate the GeneMamba’s performance to distinguish between positive pairs and negative pairs, the reference data is collected in the Gene2Vec paper [11].
 And we generate the embeddings using different models for positive gene pairs and negative gene pairs, then we measure the embeddings similarity of each pair.
@@
 To evaluate the performance of our proposed approach, we compare it with several state-of-the-art baseline methods. Below, we briefly describe each method:
-  • GeneFormer [theodoris2023transfer]: A transformer-based model specifically designed for single-cell analysis.
+  • GeneFormer [35]: A transformer-based model specifically designed for single-cell analysis.
 It uses a gene-token representation to capture complex gene-gene relationships and downstream cellular heterogeneity.
-  • scGPT [cui2024scgpt]: A generative pre-trained transformer adapted for single-cell data analysis.
+  • scGPT [5]: A generative pre-trained transformer adapted for single-cell data analysis.
 scGPT utilizes a pretraining-finetuning paradigm to handle various single-cell tasks, including cell-type classification and clustering.
-  • scFoundation [hao2024large]: A foundational model for single-cell data analysis that incorporates self-supervised learning techniques to generate robust embeddings for single-cell profiles.
+  • scFoundation [17]: A foundational model for single-cell data analysis that incorporates self-supervised learning techniques to generate robust embeddings for single-cell profiles.
 It excels in tasks requiring integration across batches and datasets.
-  • scBERT [yang2022scbert]: A BERT-like architecture adapted for single-cell data.
+  • scBERT [42]: A BERT-like architecture adapted for single-cell data.
 scBERT leverages bidirectional contextual embeddings to model complex interactions between genes within cells, making it suitable for both classification and clustering tasks.
-  • Harmony [korsunsky2019fast]: A batch-effect correction method that aligns single-cell datasets from different conditions or experiments.
+  • Harmony [20]: A batch-effect correction method that aligns single-cell datasets from different conditions or experiments.
 This method is often used as a preprocessing step to ensure integration and compatibility across datasets before applying downstream machine learning methods.
@@
 Appendix D Perturbation Analysis
-To more comprehensively evaluate the generalizability and biological relevance of GeneMamba, we conducted a perturbation prediction study using protocols established by scGPT [cui2024scgpt].
+To more comprehensively evaluate the generalizability and biological relevance of GeneMamba, we conducted a perturbation prediction study using protocols established by scGPT [5].
 Perturbation studies are critical for assessing a model’s ability to simulate cellular responses to genetic interventions—a fundamental requirement in functional genomics and drug discovery.
-We employed the Perturb-seq dataset from Norman [norman2019exploring], which include two-gene combinatorial perturbations.
+We employed the Perturb-seq dataset from Norman [24], which include two-gene combinatorial perturbations.
 Specifically, we compared Mamba-based GEARS embeddings with the original GEARS model (baseline), following the benchmarking criteria previously proposed.
@@
 Genes falling within the same interval are represented identically.
-This method, however, often leads to significant information loss, as most genes with low expression levels cluster within the same bin (e.g., the [0,1) bin [boiarsky2023deep]), masking subtle but biologically relevant differences.
+This method, however, often leads to significant information loss, as most genes with low expression levels cluster within the same bin (e.g., the [0,1) bin [2]), masking subtle but biologically relevant differences.
 Adaptive Binning: Adaptive binning, employed by scGPT, dynamically adjusts bin sizes such that each interval represents an equal proportion of expressed genes within a single cell.
 While this approach aims to better capture the distribution of gene expression, it still faces challenges related to information loss, particularly for genes with extreme expression values or biological significance.
-Value Projection Strategy: The value projection strategy, first introduced by TOSICA [chen2023transformer], has since been adopted by scCLIP [xiong2023scclip], SpaFormer [wen2023single], CellLM [zhao2023large], and CellPLM [wen2023cellplm].
+Value Projection Strategy: The value projection strategy, first introduced by TOSICA [3], has since been adopted by scCLIP [40], SpaFormer [38], CellLM [46], and CellPLM [37].
 In this approach, continuous gene expression values are directly projected into high-dimensional embeddings through linear or non-linear transformations.
@@
 Additionally, continuous embeddings may be more sensitive to noise and batch effects, necessitating robust preprocessing steps.
-Rank-Based Strategy: The rank-based strategy was first introduced in iSEEEK [shen2022universal] for single-cell transformers.
+Rank-Based Strategy: The rank-based strategy was first introduced in iSEEEK [32] for single-cell transformers.
 This approach ranks gene expression values within each cell and retains only the top-ranked genes.
@@
 E.2 Model Architectures for Single-Cell Analysis
-Recent advancements in model architectures have addressed the computational challenges of transformer-based models in single-cell RNA sequencing (scRNA-seq) analysis, with notable innovations including scTCA [scTCA], a hybrid Transformer-CNN architecture designed for imputation and denoising of single-cell read counts to enhance downstream analyses; scHyena [scHyena], a foundation model for full-length scRNA-seq analysis in brain tissues that uses a novel transformer architecture to process data without information loss, improving tasks like cell type classification and data imputation; xTrimoGene [gong2024xtrimogene], an efficient and scalable representation learner employing an asymmetric encoder-decoder transformer architecture to reduce computational requirements while maintaining high accuracy in tasks such as cell type annotation and drug combination prediction; scTransSort [scTransSort], a transformer-based model for intelligent cell type annotation that leverages self-attention mechanisms to extract features from scRNA-seq data, achieving high accuracy and performance; scGraphformer [scGraphformer], which employs transformer-based graph neural networks to provide accurate and scalable cell type annotations, uncovering cellular heterogeneity and interactions in scRNA-seq data; scGFT [scGFT], a train-free, cell-centric generative model adept at synthesizing scRNA-seq data to enhance data augmentation and analysis; White-Box Diffusion Transformer[WhiteBoxDiffusionTransformer], a hybrid model combining diffusion models and white-box transformers to generate synthetic and biologically plausible scRNA-seq data, improving training efficiency and resource utilization; STGRNS [STGRNS], an interpretable transformer-based method for inferring gene regulatory networks from scRNA-seq data to enhance understanding of gene interactions; scRDiT [scRDiT], a generative approach using diffusion transformers to create virtual scRNA-seq data, facilitating the study of gene expression dynamics; and TransformerST [TransformerST], an unsupervised model based on the transformer architecture for super-resolution in spatial transcriptomics, improving spatial gene expression analysis.
+Recent advancements in model architectures have addressed the computational challenges of transformer-based models in single-cell RNA sequencing (scRNA-seq) analysis, with notable innovations including scTCA [44], a hybrid Transformer-CNN architecture designed for imputation and denoising of single-cell read counts to enhance downstream analyses; scHyena [26], a foundation model for full-length scRNA-seq analysis in brain tissues that uses a novel transformer architecture to process data without information loss, improving tasks like cell type classification and data imputation; xTrimoGene [14], an efficient and scalable representation learner employing an asymmetric encoder-decoder transformer architecture to reduce computational requirements while maintaining high accuracy in tasks such as cell type annotation and drug combination prediction; scTransSort [18], a transformer-based model for intelligent cell type annotation that leverages self-attention mechanisms to extract features from scRNA-seq data, achieving high accuracy and performance; scGraphformer [12], which employs transformer-based graph neural networks to provide accurate and scalable cell type annotations, uncovering cellular heterogeneity and interactions in scRNA-seq data; scGFT [25], a train-free, cell-centric generative model adept at synthesizing scRNA-seq data to enhance data augmentation and analysis; White-Box Diffusion Transformer[6], a hybrid model combining diffusion models and white-box transformers to generate synthetic and biologically plausible scRNA-seq data, improving training efficiency and resource utilization; STGRNS [41], an interpretable transformer-based method for inferring gene regulatory networks from scRNA-seq data to enhance understanding of gene interactions; scRDiT [10], a generative approach using diffusion transformers to create virtual scRNA-seq data, facilitating the study of gene expression dynamics; and TransformerST [45], an unsupervised model based on the transformer architecture for super-resolution in spatial transcriptomics, improving spatial gene expression analysis.
 These models collectively advance the field by addressing key challenges in scRNA-seq data processing and interpretation.
-Traditional single-cell large language models (LLMs) are predominantly based on transformer architectures [boiarsky2023deep, white2024comprehensive].
+Traditional single-cell large language models (LLMs) are predominantly based on transformer architectures [2, 39].
 While transformers have demonstrated exceptional performance in various single-cell transcriptomics tasks, their high computational and memory requirements make training large models time-intensive and resource-heavy.
@@
 To alleviate these challenges, some transformer-based models have incorporated optimizations to improve efficiency.
-For example, scBERT integrates Performer [choromanski2020rethinking] to approximate attention calculations, while scGPT employs Flash Attention [dao2022flashattention] to accelerate training by reducing memory overhead.
+For example, scBERT integrates Performer [4] to approximate attention calculations, while scGPT employs Flash Attention [7] to accelerate training by reducing memory overhead.
 Despite these advancements, the overall time required to train such models remains substantial, especially for datasets with millions of cells and thousands of genes.
@@
 Note: Results for scBERT and scMulan were obtained by fine-tuning their publicly released pretrained models on our standardized downstream annotation datasets, using the same MLP head, learning rate, and stratified train/test split as described in Appendix B.2.
+NeurIPS Paper Checklist
+The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
+Do not remove the checklist: The papers not including the checklist will be desk rejected.
+The checklist should follow the references and follow the (optional) supplemental material.
+The checklist does NOT count towards the page limit.
+Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
+  • You should answer [Yes] , [No] , or [N/A] .
+  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
+  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
+The checklist answers are an integral part of your paper submission.
+They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
+You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
+The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
+While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
+In general, answering "[No] " or "[N/A] " is not grounds for rejection.
+While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
+All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
+If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
+IMPORTANT, please:
+  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
+  • Keep the checklist subsection headings, questions/answers and guidelines below.
+  • Do not modify the questions and only use the provided macros for your answers.
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We discussed the background scope and introduced our contributions in abstract and introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We provided the discussion about the limitations at the end of conclusion.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We provided assumptions and proofs to our theoretical results in Method and Appendix sections.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We included the link to the code for reproducibility.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We included the link to the code and dataset.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We included the link to the code for reproducibility.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We conduct our experiments in five random seeds and report the mean value in the performance table.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We provided the information of computing resources in the appendix.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Our research doesn’t have a concern about the ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: The discussion about the broader impacts is provided in the conclusion section.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We produce the data by ourselves and the original papers we refer to are cited. The baseline model we used are also cited.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We provided the details about our new proposed method in the paper.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.